### v1

# A reconstructed discontinuous approximation for distributed elliptic control problems

Abstract.
In this paper, we present and analyze an internal penalty discontinuous Galerkin method for the distributed elliptic optimal control problems. It is based on a reconstructed discontinuous approximation which admits arbitrarily high-order approximation space with only one unknown per element. Applying this method, we develop a proper discretization scheme that approximates the state and adjoint variables in the approximation space. Our main contributions are twofold: (1) the derivation of both a priori and a posteriori error estimates of the $L^{2}$-norm and the energy norms, and (2) the implementation of an efficiently solvable discrete system, which is solved via a linearly convergent projected gradient descent method. Numerical experiments are provided to verify the convergence order in a priori estimate and the efficiency of a posteriori error estimate.
keywords: distributed elliptic control; discontinuous Galerkin method; reconstructed discontinuous approximation
1. Introduction
Optimal control problems governed by partial differential equations serve as a vital bridge between advanced mathematical theory and real-world engineering. Their critical importance across applied sciences has motivated decades of research, leading to significant advancements in both the underlying theory and the numerical techniques devised to tackle them. We refer to [23] for a mathematical theory of optimal control problems and [25, 28] for some numerical methods. Among all these methods, the finite element method is extensively used to numerically solve this class of problems due to its high accuracy and adaptability.
The numerical analysis based on the finite element method can be traced back to 1970s with the works in [12, 14]. Thereafter, research on many types of control problems has advanced significantly over the past few decades. These developments range from comprehensive a priori and a posteriori error analysis to the creation of efficient error indicators for adaptive finite element methods. A description of the methodology can be found in [18]. More specifically, for Dirichlet boundary control, we refer to [10, 13, 15]. For Neumann boundary control, we refer to [14, 7, 8, 6]. For distributed control, we refer to [9, 4, 5, 22] and references therein. Variational discretization approaches for the control constrained problems are proposed in [17, 16]. In the context of adaptive finite element methods, we refer to [16, 15, 22] and references therein. The monograph [24] constructed a framework of using multi-adaptive meshes in developing efficient numerical algorithms for optimal control with error estimates.
In recent years, discontinuous Galerkin (DG) finite element methods have been widely investigated; we refer to [2, 11, 26] for some monographs. Beyond the basic feature of allowing inter‑element discontinuities, DG methods are distinguished by a combination of traits: inherent local conservation, compact stencils and high‑order accuracy with inherent parallel scalability. These attributes make them a compelling choice for a wide range of problems. On the other hand, DG methods have relatively lower computational efficiency compared to continuous elements because of the significant increase in the number of degrees of freedom [19]. This shortcoming is particularly pronounced for high‑order schemes.
To address this issue and reduce the number of degrees of freedom, the authors introduced a reconstructed discontinuous approximation (RDA) method in [21]. This method constructs a high-order approximation space requiring only one unknown per element. The core step involves solving a local least-squares problem over a patch of neighboring elements. The resulting space forms a small subspace of the standard discontinuous piecewise polynomial space, which is linearly isomorphic to the discontinuous piecewise constant space. Consequently, the number of degrees of freedom is substantially reduced to match the number of mesh elements. Moreover, it has been proved that optimal approximation accuracy can be achieved by appropriately selecting the size of the element patch. This approach has been successfully utilized to solve a series of practical problems, as documented in [21, 20].
In this article, we focus on distributed optimal control problems governed by second-order elliptic equations. Such models are ubiquitous in physics and engineering, as they describe equilibrium states in diffusion and potential flow phenomena. A typical approach to their discretization relies on classical Lagrange finite elements. For instance, [22] adopted this method and provided a priori and recovery type a posteriori estimates of the modeling error. In addition, [4, 5] studied problems with pointwise state constraints on non-convex domains, proving a priori error estimates using both continuous and discontinuous finite elements. A theoretical advancement concerns adaptive methods under pointwise control constraints. Notably, [16] provided a rigorous convergence analysis for the adaptive finite element method under variational control discretization. Furthermore, the same work established the optimality of the adaptive algorithm with respect to approximations of the state and adjoint variables.
This paper applies the aforementioned RDA method to this class of problems, which feature strongly convex cost functionals. The numerical formulation employs the approximation space for the state and adjoint variables, paired with a piecewise constant discretization for the control variable. We provide a well-defined discretization and prove both a priori and a posteriori error estimates of the $L^{2}$-norm and the energy norms. The discrete optimality conditions are solved using a projected gradient descent method, which is proved to have linear convergence rate. Each computational loop involves the sequential computation of the state and the adjoint equation via the RDA method described above, and updating the control variable using the solutions to these two equations. The performance and theoretical properties of this method are then validated through a series of numerical examples.
The rest of the article is organized as follows. In Section 2, we present the model problem and outline the basic structure of the RDA method. We put forward the discrete system and an iterative method to solve it using projected gradient descent. In Section 3, we present a primary result for a priori error estimate and a super-convergence result for the box constraint case with added regularity on the control variable. In Section 4, we prove a general a posteriori error estimate and a sharp estimate for the box constraint case. In Section 5, we present some numerical examples to confirm the theoretical results. Finally, we summarize the article in Section 6.
2. Model Problem and the RDA Method
In this section, we will describe the fundamental formulations of the model problem and the numerical method. In Section 2.1, we introduce the model problem. In Section 2.2 and 2.3, we construct the finite element space and illustrate the RDA method. In Section 2.4, a discrete system is designed to approximate the model problem. In Section 2.5, a linearly convergent iterative method is provided to solve the discrete system.
2.1. Model Problem
Let $\Omega_{u}\subset\mathbb{R}^{m}$ and $\Omega\subset\mathbb{R}^{d}(m,d=2,3)$ be convex polygonal or polyhedral domains. Take the control space $U=L^{2}(\Omega_{u}),$ the state space $V=H^{2}(\Omega)$ and the admissible set $U_{ad}\subset U$ which is closed and convex. Consider the following optimal control problem:
(2.1) $\min_{u\in U_{ad}}g(y)+j(u),$
$\mathrm{s.t.}$ $-\nabla\cdot(A\nabla y)=f+Bu\quad\mathrm{in}\ \Omega,$
$y=\phi\quad\mathrm{on}\ \partial\Omega,$
[…]
$\lambda\|\xi\|^{2}\leq\xi^{T}A(x)\xi\leq\Lambda\|\xi\|^{2}\quad\forall\xi\in\mathbb{R}^{d},\ x\in\Omega.$
In addition, we assume $f\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega),B:L^{2}(\Omega_{u})\rightarrow L^{2}(\Omega)$ is a bounded linear operator.
The state equation of problem (2.1) can be divided into two independent parts as
$\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{1})=f+Bu\quad\mathrm{in}\ \Omega,\\ &y_{1}=0\quad\mathrm{on}\ \partial\Omega,\end{aligned}\right.\quad\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{2})=0\quad\mathrm{in}\ \Omega,\\ &y_{2}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
Notice that $y_{2}$ does not rely on $u$, we denote ${g_{1}}(y_{1}):=g(y_{1}+y_{2})$. Define a bilinear form $a:H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)\rightarrow\mathbb{R}$ as
$a(y_{1},w):=(A\nabla y_{1},\nabla w)_{L^{2}(\Omega)}.$
Denote the optimal solution to (2.1) by $(y^{\ast},u^{\ast})$, then we have the state equation
(2.2) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*}\quad\mathrm{in}\ \Omega,\\ &y^{\ast}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
and the relation $y_{1}^{\ast}=y^{\ast}-y_{2}$. Define the adjoint variable $p^{\ast}$ such that
(2.3) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast})\quad\mathrm{in}\ \Omega,\quad\;\\ &p^{\ast}=0\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
Consider the Lagrange function
$L(y_{1}(u),u):={g_{1}}(y_{1}(u))+j(u)-a(y_{1}(u),p^{\ast})+(f+Bu,p^{\ast})\quad\forall u\in U_{ad}.$
Here we use (,)($\cdot\,$$\cdot\,$) to denote the $L^{2}$ inner-product $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Notice that $L(y_{1}(u),u)=g_{1}(y_{1}(u))+j(u),\forall u\in U_{ad},$ we have
$L(y_{1}(u^{\ast}+t(v-u^{\ast})),u^{\ast}+t(v-u^{\ast}))\geq L(y_{1}^{\ast},u^{\ast}),\quad\forall v\in U_{ad},\ t\in[0,1].$
Differentiate the above inequality at $t=0$ in the direction $v-u^{*}.$ We have
$(g_{1}^{\prime}(y_{1}^{*})y_{1}^{\prime}(u^{*})+j^{\prime}(u^{*}))(v-u^{*})-a(y_{1}^{\prime}(u^{*})(v-u^{*}),p^{*})+(B(v-u^{*}),p^{*})\geq 0.$
Since
$a(w,p^{*})=(g^{\prime}(y^{\ast}),w)=(g_{1}^{\prime}(y_{1}^{\ast}),w),\quad\forall w\in H_{0}^{1}(\Omega),$
we take $w=y_{1}^{\prime}(u^{*})(v-u^{*})\in H_{0}^{1}(\Omega)$ and derive
$(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
So far we have derived the optimality conditions of problem (2.1) as follows.
(2.4) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*},\quad y^{\ast}|_{\partial\Omega}=\phi,\\ &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast}),\quad p^{\ast}|_{\partial\Omega}=0,\\ &(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.\end{aligned}\right$.
It is well-known from [23] that there exists a unique solution ($y^{*},u^{*}$) to problem (2.1), and the optimality condition (2.4) is a necessary and sufficient condition of the exact solution ($y^{*},u^{*}$).
Assumption 1.
[…]
$\alpha\|u-\tilde{u}\|_{L^{2}}^{2}\leq(j^{\prime}(u)-j^{\prime}(\tilde{u}),u-\tilde{u})_{L^{2}},\quad\beta\|y-\tilde{y}\|_{L^{2}}^{2}\leq(g^{\prime}(y)-g^{\prime}(\tilde{y}),y-\tilde{y})_{L^{2}},$
$\|j^{\prime}(u)-j^{\prime}(\tilde{u})\|_{L^{2}}\leq C\|u-\tilde{u}\|_{L^{2}},\quad\|g^{\prime}(y)-g^{\prime}(\tilde{y})\|_{L^{2}}\leq C\|y-\tilde{y}\|_{L^{2}},$
for $\forall u,\tilde{u}\in U_{ad},y,\tilde{y}\in Y$. Here $\alpha,\beta,C$ are positive constants.
Remark 1.
Here we denote $\|u-\tilde{u}\|_{L^{2}}:=\|u-\tilde{u}\|_{L^{2}(\Omega_{u})}$ and $\|y-\tilde{y}\|_{L^{2}}:=\|y-\tilde{y}\|_{L^{2}(\Omega)}.$ We will use these notations in the following part if there is no confusion.
Assumption 2.
The functionals $j(u)$ and $g(y)$ can be written in the form
$j(u)=\int_{\Omega_{u}}\tilde{j}(u)\mathrm{d}x,\quad g(y)=\int_{\Omega}\tilde{g}(y)\mathrm{d}x,$
which means that
$(j^{\prime}(u),v)=\int_{\Omega_{u}}\tilde{j}^{\prime}(u)v\mathrm{d}x,\quad(g^{\prime}(y),w)=\int_{\Omega}\tilde{g}^{\prime}(y)w\mathrm{d}x,$
for $\forall u,v\in U,y,w\in Y.$
[…]
2.2. The Reconstruction Operator
We are going to construct the finite element space. Let $\mathcal{T}_{h}$ be a quasi-uniform triangulation over $\Omega\subset\mathbb{R}^{d}$ with mesh size $h$. Let $\mathcal{E}_{h}^{I}$ and $\mathcal{E}_{h}^{B}$ be the set of all d1d-1 dimensional faces in the interior of $\Omega$ in $\mathcal{T}_{h}$, and faces on the boundary $\partial\Omega$ in $\mathcal{T}_{h}$ respectively. The set of all faces is denoted by $\mathcal{E}_{h}:=\mathcal{E}_{h}^{I}\cup\mathcal{E}_{h}^{B}.$ Now we will briefly introduce the reconstruction operator $\mathcal{R}^{m}$ proposed in [21]. The construction of $\mathcal{R}^{m}$ includes three steps in total.
Step 1. Given $K\in\mathcal{T}_{h}$, we want to construct an element patch $S(K)$, which consists of $K$ itself and some of its surrounding elements. The construction of $S(K)$ is conducted in a recursive way. We begin by setting $S_{0}(K)=\{K\}$, and define $S_{t}(K)$ recursively:
$S_{t}(K)=\bigcup\limits_{K^{\prime}\in S_{t-1}(K)}N(K^{\prime}),\quad t=0,1,\cdots$
where
$N(K):=\{K^{\prime}\in\mathcal{T}_{h}\mid\exists e\in\mathcal{E}_{h}\ \mathrm{s.t.}\ e\subset\partial{K^{\prime}}\cap\partial{K}\}.$
The recursion stops once $t$ meets the condition that the cardinality $S_{t}(K)\geq\#S$, which is a given constant independent of $K$. Then we define the patch $S(K):=S_{t}(K).$ Generally, the threshold #S\#S can be chosen as $\frac{d+1}{2}\mathrm{dim}(\mathbb{P}_{m}(K))$ by numerical observations.
Step 2. For each $K\in\mathcal{T}_{h}$, we will solve a local least square problem on $S(K).$ Let $\boldsymbol{x}_{K}$ be the barycenter of $K$ and $I(K):=\{\boldsymbol{x}_{K^{\prime}}\in\Omega\mid K^{\prime}\in S(K)\}$ be the set of barycenters of elements in $S(K)$. Let $V_{h}^{0}$ be the piecewise constant space on $\mathcal{T}_{h}$,
$V_{h}^{0}:=\{w_{h}\in L^{2}(\Omega)\mid w_{h}|_{K}\in\mathbb{P}_{0}(K),\>\forall K\in\mathcal{T}_{h}\}.$
Given a piecewise constant function $w_{h}\in V_{h}^{0}$, for each $K\in\mathcal{T}_{h}$, we seek a polynomial of degree no more than $m$ by the following constrained local least squares problem:
(2.5) $\underset{p\in\mathbb{P}_{m}(S(K))}{\operatorname*{\min}}\sum_{\boldsymbol{x}\in I(K)}(p(\boldsymbol{x})-w_{h}(\boldsymbol{x}))^{2},$
$\mathrm{s.t.}$ $p(\boldsymbol{x}_{K})=w_{h}(\boldsymbol{x}_{K}).$
The existence and uniqueness of the solution to this problem is detailed in [21]. We denote this solution as $w_{h,S(K)}\in\mathbb{P}_{m}(S(K))$.
Step 3. Given $K\in\mathcal{T}_{h}$, notice that $w_{h,S(K)}$ linearly depends on the function $w_{h}$, we can then define a linear operator $\mathcal{R}_{K}^{m}:V_{h}^{0}\longrightarrow\mathbb{P}_{m}(S(K))$ mapping $w_{h}$ to $w_{h,S(K)}$. Finally, we define the global reconstruction operator $\mathcal{R}^{m}$ as
$\mathcal{R}^{m}:\>V_{h}^{0}\longrightarrow V_{h}^{m},$
$\mathrm{s.t.}$ $(\mathcal{R}^{m}w_{h})|_{K}=(\mathcal{R}_{K}^{m}w_{h})|_{K},\quad\forall K\in\mathcal{T}_{h}.$
Here $V_{h}^{m}:=\mathcal{R}^{m}V_{h}^{0}$ is the image space. $(\mathcal{R}^{m}w_{h})|_{K}$ is the restriction of $\mathcal{R}_{K}^{m}w_{h}$ on $K\in\mathcal{T}_{h}$. Any piecewise constant function $w_{h}$ is mapped to a piecewise $m$-th degree polynomial $\mathcal{R}^{m}w_{h}$. Moreover, it can be shown that $\mathcal{R}^{m}$ is a linear isomorphism from $V_{h}^{0}$ to $V_{h}^{m}$, i.e. $\mathrm{dim}(V_{h}^{m})=\mathrm{dim}(V_{h}^{0}).$
2.3. The Reconstructed Discontinuous Approximation Method
Here we present the numerical method for the second-order elliptic problem based on the interior penalty DG method and the reconstructed space $V_{h}^{m}$.
We first introduce some notations commonly used in the DG framework. Let $e\in\mathcal{E}_{h}^{I}$ be any interior face shared by $K^{+},K^{-}\in\mathcal{T}_{h}$ with the unit outward normal vectors $\mathbf{n}^{+},\mathbf{n}^{-}$ along $e$, respectively. For any piecewise smooth scalar-valued function $v$ and vector-valued function $\boldsymbol{\tau}$, the jump operator $\cdot\,$] and the average operator $\{\,\cdot\,\}$ are defined as
$v]|_{e}:=v^{+}|_{e}\mathbf{n}^{+}+v^{-}|_{e}\mathbf{n}^{-},\quad\{v\}|_{e}:=\frac{1}{2}(v^{+}|_{e}+v^{-}|_{e}),$
$\boldsymbol{\tau}]|_{e}:=\boldsymbol{\tau}^{+}|_{e}\cdot\mathbf{n}^{+}+\boldsymbol{\tau}^{-}|_{e}\cdot\mathbf{n}^{-},\quad\{\boldsymbol{\tau}\}|_{e}:=\frac{1}{2}(\boldsymbol{\tau}^{+}|_{e}+\boldsymbol{\tau}^{-}|_{e}),$
where $v^{\pm}:=v|_{K^{\pm}},\boldsymbol{\tau}^{\pm}:=\boldsymbol{\tau}|_{K^{\pm}}.$ For any boundary face $e\in\mathcal{E}_{h}^{B}$, with the unit outward normal $\mathbf{n}$ along $e,$ the above two operators are modified as
$v]|_{e}:=v|_{e}\mathbf{n},\ \{v\}|_{e}:=v|_{e},\quad[\boldsymbol{\tau}]|_{e}:=\boldsymbol{\tau}|_{e}\cdot\mathbf{n},\ \{\boldsymbol{\tau}\}|_{e}:=\boldsymbol{\tau}|_{e}.$
Consider the problem
(2.6) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y)=F\quad\mathrm{in}\ \Omega,\\ &y=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
Here $F\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega)$ is given. We seek the numerical solution $y_{h}^{m}\in V_{h}^{m}$ such that
(2.7) $a_{h}(y_{h}^{m},w_{h}^{m})=l_{h}(w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
where $a_{h}:V_{h}^{m}\times V_{h}^{m}\rightarrow\mathbb{R},l_{h}:V_{h}^{m}\rightarrow\mathbb{R}$ are given by
$a_{h}(y_{h}^{m},w_{h}^{m})$= $\sum_{K\in\mathcal{T}_{h}}\int_{K}A\nabla y_{h}^{m}\cdot\nabla w_{h}^{m}\mathrm{d}x-\sum_{e\in\mathcal{E}_{h}}\int_{e}\{A\nabla y_{h}^{m}\}\cdot[w_{h}^{m}]\mathrm{d}s$
- $\sum_{e\in\mathcal{E}_{h}}\int_{e}\{A\nabla w_{h}^{m}\}\cdot[y_{h}^{m}]\mathrm{d}s+\mu\sum_{e\in\mathcal{E}_{h}}h_{e}^{-1}\int_{e}[y_{h}^{m}]\cdot[w_{h}^{m}]\mathrm{d}s,$
$l_{h}(w_{h}^{m}):=\sum_{K\in\mathcal{T}_{h}}\int_{K}Fw_{h}^{m}\mathrm{d}x-\sum_{e\in\mathcal{E}_{h}^{B}}\int_{e}\{A\nabla w_{h}^{m}\}\cdot\mathbf{n}_{e}\phi\mathrm{d}s+\mu\sum_{e\in\mathcal{E}_{h}^{B}}h_{e}^{-1}\int_{e}[w_{h}^{m}]\cdot\mathbf{n}_{e}\phi\mathrm{d}s.$
Here $\mu>0$ is the penalty parameter. Denote
$H^{s}(\mathcal{T}_{h}):=\{w_{h}\in L^{2}(\Omega)\mid w_{h}|_{K}\in H^{s}(K),\forall K\in\mathcal{T}_{h}\},$
and $\widetilde{h_{e}}=h_{e}/\mu.$ We define the following DG-norms
$\|w_{h}\|_{\mathrm{DG}}^{2}:=\sum_{K\in\mathcal{T}_{h}}\|\nabla w_{h}\|_{L^{2}(K)}^{2}+\sum_{e\in\mathcal{E}_{h}}\widetilde{h_{e}}^{-1}\|[w_{h}]\|_{L^{2}(e)}^{2},\quad\forall w_{h}\in H^{1}(\mathcal{T}_{h}),$
$\interleave w_{h}\interleave_{\mathrm{DG}}^{2}:=\|w_{h}\|_{\mathrm{DG}}^{2}+\sum_{e\in\mathcal{E}_{h}}\widetilde{h_{e}}\|\{\nabla w_{h}\}\|_{L^{2}(e)}^{2},\quad\forall w_{h}\in H^{2}(\mathcal{T}_{h}).$
By the inverse estimate and the trace inequality, we know that both norms are equivalent restricted on finite dimensional space $V_{h}^{m}$, i.e.
$\|w_{h}^{m}\|_{\mathrm{DG}}\leq\interleave w_{h}^{m}\interleave_{\mathrm{DG}}\leq C\|w_{h}^{m}\|_{\mathrm{DG}},\quad\forall w_{h}^{m}\in V_{h}^{m}.$
To enlarge the domain of the bilinear form $a_{h}$, we define $V_{h}:=V_{h}^{m}+H^{2}(\Omega)$. The relationship between the energy norms and the $L^{2}$-norm in $V_{h}$ is given in [1]:
(2.8) $\|w_{h}\|_{L^{2}}\leq C\|w_{h}\|_{\mathrm{DG}}\leq C\interleave w_{h}\interleave_{\mathrm{DG}},\quad\forall w_{h}\in V_{h}.$
This constant $C$ only depends on $\Omega$ and the lower bound $K_{0}$ of all angles of all triangles in $\mathcal{T}_{h}$ and the grade constant $K_{1}$ of $\mathcal{T}_{h}$ defined as follows.
$K_{1}:=\sup\limits_{K\in\mathcal{T}_{h}}\sup\limits_{e\in\mathcal{E}_{K}}\frac{h_{K}}{h_{e}}.$
Here $\mathcal{E}_{K}=\{e\in\mathcal{E}_{h}\mid e\subset\partial K\}.$ So it does not change under uniform refinement of $\mathcal{T}_{h}.$
We can extend the domain of $a_{h},l_{h}$ to $V_{h}$ without confusion, namely
$a_{h}:V_{h}\times V_{h}\rightarrow\mathbb{R},\quad l_{h}:V_{h}\rightarrow\mathbb{R}.$
Subsequently, we can easily derive the stability of $a_{h}$ on $(H^{2}(\Omega)\cap H_{0}^{1}(\Omega),\|\cdot\|_{\mathrm{DG}})$
$a_{h}(w_{h},w_{h})\geq c\|w_{h}\|_{\mathrm{DG}}^{2},\quad\forall w_{h}\in H^{2}(\Omega)\cap H_{0}^{1}(\Omega).$
This is due to the terms of integration on every edge vanishes for $w_{h}\in H^{2}(\Omega)\cap H_{0}^{1}(\Omega)$. Naturally, we have the stability of $a_{h}$ on $(H^{2}(\Omega)\cap H_{0}^{1}(\Omega),\|\cdot\|_{L^{2}})$ by (2.8). In addition, for the exact solution $y$ to (2.6), we know by direct calculation that
$a_{h}(y,w_{h}^{m})=l_{h}(w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
which leads to the consistency of $a_{h}$
(2.9) $a_{h}(y,w_{h})=l_{h}(w_{h}),\quad\forall w_{h}\in V_{h}.$
Now we present some results for the boundedness and stability of $a_{h}$ on $(V_{h}^{m},\interleave\cdot\interleave_{\mathrm{DG}})$ and a priori and a posteriori error estimates for the reconstructed discontinuous approximation method. Their proofs refer to [21, 27].
Lemma 1.
Let $a_{h}$ be defined with sufficiently large $\mu$, then there exist constants $C,c$ such that
$|a_{h}(y_{h},w_{h})|\leq C\interleave y_{h}\interleave_{\mathrm{DG}}\interleave w_{h}\interleave_{\mathrm{DG}},\quad\forall y_{h},w_{h}\in V_{h},$
$a_{h}(w_{h}^{m},w_{h}^{m})\geq c\interleave w_{h}^{m}\interleave_{\mathrm{DG}}^{2},\quad\forall w_{h}^{m}\in V_{h}^{m}.$
Theorem 1.
Let $y\in H^{m+1}(\Omega)$ be the exact solution to (2.6), and $y_{h}^{m}\in V_{h}^{m}$ be the numerical solution to (2.7). Take $\mu$ sufficiently large, then there exists a constant $C=C(y)$ such that
$\interleave y-y_{h}^{m}\interleave_{\mathrm{DG}}\leq Ch^{m}\|y\|_{H^{m+1}},$
and
$\|y-y_{h}^{m}\|_{L^{2}}\leq Ch^{m+1}\|y\|_{H^{m+1}}.$
Theorem 2.
Let $y\in H^{2}(\Omega)$ be the exact solution to (2.6), and $y_{h}^{m}\in V_{h}^{m}$ be the numerical solution to (2.7). Take $\mu$ sufficiently large, then there exists a constant $C$ such that
$\|y-y_{h}^{m}\|_{L^{2}}\leq C(\eta_{1}(y_{h}^{m})+\eta_{2}(y_{h}^{m})+\eta_{3}(y_{h}^{m})),$
where
(2.10) $\begin{array}[]{c}\displaystyle\eta_{1}^{2}(y_{h}^{m})=\sum_{K\in\mathcal{T}_{h}}\widetilde{h_{K}}^{4}\|F+\nabla\cdot(A\nabla y_{h}^{m})\|_{L^{2}(K)}^{2},\\ \displaystyle\eta_{2}^{2}(y_{h}^{m})=\sum_{e\in\mathcal{E}_{h}}\widetilde{h_{e}}\|[y_{h}^{m}]\|_{L^{2}(e)}^{2},\quad\eta_{3}^{2}(y_{h}^{m})=\sum_{e\in\mathcal{E}_{h}^{I}}\widetilde{h_{e}}^{3}\|[A\nabla y_{h}^{m}]\|_{L^{2}(e)}^{2}.\end{array}$
Here $\widetilde{h_{K}}=h_{K}/\mu.$
2.4. Discrete System.
[…]
By the consistency (2.9), we have
(2.11) $\left\{\begin{aligned} &a_{h}(y^{\ast},w_{h})=l_{h}[f+Bu^{*},\phi](w_{h}),\quad\forall w_{h}\in V_{h},\quad\ \ \;\\ &a_{h}(q_{h},p^{\ast})=l_{h}[g^{\prime}(y^{\ast}),0](q_{h}),\quad\forall q_{h}\in V_{h},\\ &(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.\end{aligned}\right$.
We define the discretization of (2.11) as follows.
(2.12) $\left\{\begin{aligned} &a_{h}(y_{h}^{m\ast},w_{h}^{m})=l_{h}[f+Bu_{h}^{*},\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},\\ &a_{h}(q_{h}^{m},p_{h}^{m\ast})=l_{h}[g^{\prime}(y_{h}^{m\ast}),0](q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m},\\ &(j^{\prime}(u_{h}^{*})+B^{*}p_{h}^{m*},v_{h}-u_{h}^{*})\geq 0,\quad\forall v_{h}\in U_{adh}.\end{aligned}\right$.
Here for $w_{h}\in V_{h},$
$l_{h}[s,\phi](w_{h}):=\sum_{K\in\mathcal{T}_{h}}\int_{K}sw_{h}\mathrm{d}x-\sum_{e\in\mathcal{E}_{h}^{B}}\int_{e}\{A\nabla w_{h}\}\cdot\mathbf{n}_{e}\phi\mathrm{d}s+\mu\sum_{e\in\mathcal{E}_{h}^{B}}h_{e}^{-1}\int_{e}[w_{h}]\cdot\mathbf{n}_{e}\phi\mathrm{d}s,$
and $U_{adh}$ is a discretization of $U_{ad}$. Specifically, given a quasi-uniform mesh $\mathcal{T}_{h}^{u}$ on $\Omega_{u}$ with mesh size $h_{u},$ we take $U_{h}$ to be the piecewise constant function space on $\mathcal{T}_{h}^{u}$ and define $U_{adh}:=U_{ad}\cap U_{h}$, as in [24, 22, 29].
Remark 2.
Correspondingly, we can consider the discrete optimal control problem
(2.13) $\min_{u_{h}\in U_{adh}}g(y_{h}^{m})+j(u_{h}),$
$\mathrm{s.t.}$ $\ a_{h}(y_{h}^{m},w_{h}^{m})=l_{h}[f+Bu_{h},\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$\ y_{h}^{m}\in V_{h}^{m}.$
Take $\mu$ sufficiently large, we shall prove that the discrete system (2.12) is exactly the optimality condition of the discrete problem (2.13). Then the finite dimensionality of (2.13) implies the existence of the solution to (2.12). The strong convexity in Assumption 1 guarantees the uniqueness of the solution to (2.12).
Proof.
Let $(y_{h}^{m*},u_{h}^{*})$ be the solution to (2.13). The Lagrange function
$L(y_{h}^{m}(u_{h}),u_{h}):={g}(y_{h}^{m}(u_{h}))+j(u_{h})-(g^{\prime}(y_{h}^{m*}),y_{h}^{m}(u_{h})-F_{h}(f+Bu_{h},\phi)).$
[…]
$\left\{\begin{aligned} &-\nabla\cdot(A\nabla y)=w\quad\mathrm{in}\ \Omega,\\ &y=q\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
We have $L(y_{h}^{m}(u_{h}),u_{h})=g(y_{h}^{m}(u_{h}))+j(u_{h}),\forall u_{h}\in U_{adh}.$ So
$L(y_{h}^{m}(u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast})),u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast}))\geq L(y_{h}^{m*},u_{h}^{*}),\quad\forall v_{h}\in U_{adh},\ t\in[0,1].$
Differentiate the above inequality at $t=0$ in the direction $v_{h}-u_{h}^{*},$ we have
$(\tilde{g}^{\prime}(y_{h}^{m*})y_{h}^{m\prime}(u_{h}^{*})+\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})$
- $(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m\prime}(u_{h}^{*})(v_{h}-u_{h}^{*})-F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
Here we use Assumption 2. Then
$(\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})+(\tilde{g}^{\prime}(y_{h}^{m*}),F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
For $\forall w\in L^{2}(\Omega),$ we have
$(\tilde{g}^{\prime}(y_{h}^{m*}),F_{h}(w,0))=l_{h}[\tilde{g}^{\prime}(y_{h}^{m*}),0](F_{h}(w,0))$
$=$ $a_{h}(F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0),F_{h}(w,0))=a_{h}(F_{h}(w,0),F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0))$
$=$ $l_{h}[w,0](F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0))=(F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0),w).$
[…]
2.5. Iterative Method
It should be noted that the approximate system (2.12) can not be solved directly. Here we introduce an iterative method. Take $u_{h0}\in U_{adh}$ as an initial guess, we can then obtain a sequence {($u_{hn},y_{hn}^{m},p_{hn}^{m})\}_{n=0}^{\infty}$ by the following algorithm.
(2.14) $\left\{\begin{aligned} &a_{h}(y_{hn}^{m},w_{h}^{m})=l_{h}[f+Bu_{hn},\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},\\ &a_{h}(q_{h}^{m},p_{hn}^{m})=l_{h}[g^{\prime}(y_{hn}^{m}),0](q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m},\\ &u_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}=Pr_{U_{adh}}(u_{hn}-\rho_{n}(j^{\prime}(u_{hn})+B^{*}p_{hn}^{m})),\end{aligned}\right$.
where $Pr_{U_{adh}}:U\rightarrow U_{adh}$ is the projection operator and $\rho_{n}>0,\forall n\in\mathbb{N}.$ This is essentially a projected gradient descent method, which has linear convergence rate under strongly convex cost function. We will prove this property in detail.
Lemma 2.
Let $(u_{h},y_{h}^{m},p_{h}^{m})$ satisfy (2.12), $\{(u_{hn},y_{hn}^{m},p_{hn}^{m})\}_{n=0}^{\infty}$ be the sequence generated by (2.14). If Assumptions 1,2 hold, we can properly choose $\{\rho_{n}\}_{n=0}^{\infty}$ such that there exist constants $0<\eta<1,\lambda=\lambda(\eta)>0,$ satisfy
$\|u_{hn}-u_{h}\|_{L^{2}}+\lambda\interleave y_{hn}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}+\lambda\interleave p_{n}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}$
$\leq$ $\eta^{n}(\|u_{h0}-u_{h}\|_{L^{2}}+\lambda\interleave y_{h0}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}+\lambda\interleave p_{h0}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}).$
In fact, we can show that $(u_{hn},y_{hn}^{m},p_{hn}^{m})$ converges to $(u_{h},y_{h}^{m},p_{h}^{m})$ linearly.
Proof.
We have
$\|u_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}-u_{h}\|_{L^{2}}^{2}=$ $\|Pr_{U_{adh}}(u_{hn}-\rho_{n}(j^{\prime}(u_{hn})+B^{*}p_{hn}^{m}))$
$-Pr_{U_{adh}}(u_{h}-\rho_{n}(j^{\prime}(u_{h})+B^{*}p_{h}^{m}))\|_{L^{2}}^{2}$
$\leq$ $\|u_{hn}-u_{h}-\rho_{n}(j^{\prime}(u_{hn})-j^{\prime}(u_{h})+B^{*}p_{hn}^{m}-B^{*}p_{h}^{m})\|_{L^{2}}^{2}$
$=$ $\|u_{hn}-u_{h}\|_{L^{2}}^{2}+\rho_{n}^{2}\|j^{\prime}(u_{hn})-j^{\prime}(u_{h})+B^{*}p_{hn}^{m}-B^{*}p_{h}^{m}\|_{L^{2}}^{2}$
$-2\rho_{n}(u_{hn}-u_{h},j^{\prime}(u_{hn})-j^{\prime}(u_{h})+B^{*}p_{hn}^{m}-B^{*}p_{h}^{m})$
$\leq$ $(1-2\alpha\rho_{n}+C\rho_{n}^{2})\|u_{hn}-u_{h}\|_{L^{2}}^{2}+C\rho_{n}^{2}\|p_{hn}^{m}-p_{h}^{m}\|_{L^{2}}^{2}$
$-2\rho_{n}(u_{hn}-u_{h},B^{*}p_{hn}^{m}-B^{*}p_{h}^{m}).$
From (2.12) and (2.14) we have
(2.15) $a_{h}(y_{hn}^{m}-y_{h}^{m},w_{h}^{m})=(B(u_{hn}-u_{h}),w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$a_{h}(q_{h}^{m},p_{hn}^{m}-p_{h}^{m})=(g^{\prime}(y_{hn}^{m})-g^{\prime}(y_{h}^{m}),q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m}.$
Let $w_{h}^{m}=p_{hn}^{m}-p_{h}^{m},q_{hn}^{m}=y_{hn}^{m}-y_{h}^{m}.$ Then
$(u_{hn}-u_{h},B^{*}p_{hn}^{m}-B^{*}p_{h}^{m})=(B(u_{hn}-u_{h}),p_{hn}^{m}-p_{h}^{m})$
$=$ $(g^{\prime}(y_{hn}^{m}))-g^{\prime}(y_{h}^{m}),y_{hn}^{m}-y_{h}^{m})\geq\beta\|y_{hn}^{m}-y_{h}^{m}\|_{L^{2}}^{2}\geq 0.$
Let $w_{h}^{m}=y_{hn}^{m}-y_{h}^{m},q_{hn}^{m}=p_{hn}^{m}-p_{h}^{m}.$ By the stability of $a_{h}$ on $(V_{h}^{m},\interleave\cdot\interleave_{\mathrm{DG}})$ and (2.8), we have
$\interleave y_{hn}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}\leq C\|u_{hn}-u_{h}\|_{L^{2}},\quad\interleave p_{hn}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}\leq C\|y_{hn}^{m}-y_{h}^{m}\|_{L^{2}}.$
Consequently,
$\|u_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}-u_{h}\|_{L^{2}}^{2}\leq(1-2\alpha\rho_{n}+C\rho_{n}^{2})\|u_{hn}-u_{h}\|_{L^{2}}^{2}\leq\delta^{2}\|u_{hn}-u_{h}\|_{L^{2}}^{2},$
where $\rho_{n}$ satisfies $0\leq 1-2\alpha\rho_{n}+C\rho_{n}^{2}\leq\delta^{2}<1,\forall n\in\mathbb{N}.$ Choose $C^{\prime}$ such that
$\interleave y_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}+\interleave p_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}\leq C^{\prime}\|u_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}-u_{h}\|_{L^{2}},$
and $\lambda=(1-\delta)/(2C^{\prime}\delta)$. It can be verified that
$\|u_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}-u_{h}\|_{L^{2}}+\lambda\interleave y_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}+\lambda\interleave p_{h\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}$
$\leq$ $\frac{\delta+1}{2}(\|u_{hn}-u_{h}\|_{L^{2}}+\lambda\interleave y_{hn}^{m}-y_{h}^{m}\interleave_{\mathrm{DG}}+\lambda\interleave p_{hn}^{m}-p_{h}^{m}\interleave_{\mathrm{DG}}),$
which leads to the linear convergence rate. To complete the proof, we only need to take $\eta=(\delta+1)/2$. ∎
[…]
3. A Priori Estimates
In this section, we will establish a priori error estimate of the discrete system (2.12) with respect to the optimality condition (2.11). We will first prove a theorem in general case, and then improve the result to a sharper estimate for the box constraint case.
3.1. A General Result
To complete the foundation for the proof below, we introduce one further technical assumption, which will be employed in the following part.
Assumption 3.
Let $\Pi_{h}^{u}:U\rightarrow U_{h}$ be the orthogonal projection to $U_{h}$ under the $L^{2}$-norm, such that $(u-\Pi_{h}^{u}u,v_{h})=0,\forall v_{h}\in U_{h}.$ We shall assume that
$\Pi_{h}^{u}U_{ad}\subset{U_{adh}},\quad\mathrm{i.e.}\ \Pi_{h}^{u}v\in U_{adh},\forall v\in U_{ad}.$
It should be noted that Assumption 3 is available for the following cases:
$U_{ad}=U,\quad U_{ad}=\{v\in U\mid\gamma\leq\int_{\Omega_{u}}v\mathrm{d}x\leq\Gamma\},\quad U_{ad}=\{v\in U\mid\gamma\leq v\leq\Gamma\ \mathrm{a.e.\ in}\ \Omega_{u}\},$
where $\gamma\leq\Gamma$ are constants.
Theorem 3.
[…]
$\|u-u_{h}\|_{L^{2}}+\interleave y-y_{h}^{m}\interleave_{\mathrm{DG}}+\interleave p-p_{h}^{m}\interleave_{\mathrm{DG}}\leq C(h_{u}+h^{m}),$
and
$\|u-u_{h}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}+\|p-p_{h}^{m}\|_{L^{2}}\leq C(h_{u}+h^{m+1}).$
Proof.
We construct the intermediate states as follows.
(3.1) $a_{h}(P_{h}y,w_{h}^{m})=l_{h}[f+Bu,\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$a_{h}(q_{h}^{m},P_{h}p)=l_{h}[g^{\prime}(y),0](q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m}.$
$P_{h}y,P_{h}p\in V_{h}^{m}.$ From (2.12) and (3.1), we have
(3.2) $a_{h}(P_{h}y-y_{h}^{m},w_{h}^{m})=(B(u-u_{h}),w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$a_{h}(q_{h}^{m},P_{h}p-p_{h}^{m})=(g^{\prime}(y)-g^{\prime}(y_{h}^{m}),q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m}.$
Take $w_{h}^{m}=P_{h}p-p_{h}^{m},q_{h}^{m}=P_{h}y-y_{h}^{m}$ in (3.2) and subtract the resulting equations to find
$(B^{*}(P_{h}p-p_{h}^{m}),u-u_{h})=(g^{\prime}(y)-g^{\prime}(y_{h}^{m}),P_{h}y-y_{h}^{m}).$
This implies
(3.3) $(B^{*}(P_{h}p-p_{h}^{m}),u-u_{h})\geq\beta\|P_{h}y-y_{h}^{m}\|^{2}_{L^{2}}+(g^{\prime}(y)-g^{\prime}(P_{h}y),P_{h}y-y_{h}^{m}).$
From the inequalities in (2.11) and (2.12), we know that $(j^{\prime}(u)+B^{*}p,u_{h}-u)\geq 0$ and
$(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-u_{h})\geq(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-v_{h}),\quad\forall v_{h}\in U_{adh}.$
Thus
$(j^{\prime}(u_{h})-j^{\prime}(u)+B^{*}(p_{h}^{m}-p),u-u_{h})\geq(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-v_{h}),\quad\forall v_{h}\in U_{adh},$
which implies
(3.4) $(B^{*}(p_{h}^{m}-p),u-u_{h})\geq\alpha\|u-u_{h}\|_{L^{2}}^{2}+(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-v_{h}),\quad\forall v_{h}\in U_{adh}.$
Take $v_{h}=\Pi_{h}^{u}u$ (by Assumption 3).
Combine (3.3) and (3.4), we have
$\alpha\|u-u_{h}\|_{L^{2}}^{2}+\beta\|y_{h}^{m}-P_{h}y\|_{L^{2}}^{2}$
$\leq$ $-(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-\Pi_{h}^{u}u)-(B^{*}(p-P_{h}p),u-u_{h})+(g^{\prime}(y)-g^{\prime}(P_{h}y),y_{h}^{m}-P_{h}y)$
$=$ $-(j^{\prime}(u)+B^{*}p,u-\Pi_{h}^{u}u)-(j^{\prime}(u_{h})-j^{\prime}(u)+B^{*}(p_{h}^{m}-p),u-\Pi_{h}^{u}u)$
$-(B^{*}(p-P_{h}p),u-u_{h})+(g^{\prime}(y)-g^{\prime}(P_{h}y),y_{h}^{m}-P_{h}y)$
$=$ $-((I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p),u-\Pi_{h}^{u}u)-(j^{\prime}(u_{h})-j^{\prime}(u)+B^{*}(p_{h}^{m}-p),u-\Pi_{h}^{u}u)$
$-(B^{*}(p-P_{h}p),u-u_{h})+(g^{\prime}(y)-g^{\prime}(P_{h}y),y_{h}^{m}-P_{h}y)$
$\leq$ $C(\|(I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p)\|_{L^{2}}^{2}+\|u-\Pi_{h}^{u}u\|_{L^{2}}^{2}+\|B^{*}(p_{h}^{m}-p)\|_{L^{2}}\|u-\Pi_{h}^{u}u\|_{L^{2}}\quad\quad\ \$
$+\|B^{*}(p-P_{h}p)\|_{L^{2}}^{2}+\|y-P_{h}y\|_{L^{2}}^{2})+\frac{\alpha}{2}\|u-u_{h}\|_{L^{2}}^{2}+\frac{\beta}{2}\|y_{h}^{m}-P_{h}y\|_{L^{2}}^{2}.$
[…]
$\|P_{h}p-p_{h}^{m}\|_{L^{2}}\leq C\interleave P_{h}p-p_{h}^{m}\interleave_{\mathrm{DG}}\leq C\|y-y_{h}^{m}\|_{L^{2}}.$
Thus
$\|p-p_{h}^{m}\|_{L^{2}}=$ $\|p-P_{h}p+P_{h}p-p_{h}^{m}\|_{L^{2}}\leq C(\|y-y_{h}^{m}\|_{L^{2}}+\|p-P_{h}p\|_{L^{2}})$
$\leq$ $C(\|y-P_{h}y\|_{L^{2}}+\|y_{h}^{m}-P_{h}y\|_{L^{2}}+\|p-P_{h}p\|_{L^{2}}).$
So we have
$\alpha\|u-u_{h}\|_{L^{2}}^{2}+\beta\|y_{h}^{m}-P_{h}y\|_{L^{2}}^{2}$
$\leq$ $C(\|(I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p)\|_{L^{2}}^{2}+\|u-\Pi_{h}^{u}u\|_{L^{2}}^{2}+\|p-P_{h}p\|_{L^{2}}^{2}+\|y-P_{h}y\|_{L^{2}}^{2})$
$+\frac{\alpha}{2}\|u-u_{h}\|_{L^{2}}^{2}+\frac{3\beta}{4}\|y_{h}^{m}-P_{h}y\|_{L^{2}}^{2},$
which means that
$\|u-u_{h}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}$
$\leq$ $C(\|(I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p)\|_{L^{2}}+\|u-\Pi_{h}^{u}u\|_{L^{2}}+\|p-P_{h}p\|_{L^{2}}+\|y-P_{h}y\|_{L^{2}})$
by the triangular inequality.
Now we give upper bounds for $\interleave y-y_{h}^{m}\interleave_{\mathrm{DG}}$ and $\interleave p-p_{h}^{m}\interleave_{\mathrm{DG}}.$ We have known that
$\interleave p-p_{h}^{m}\interleave_{\mathrm{DG}}\leq C(\interleave p-P_{h}p\interleave_{\mathrm{DG}}+\|y-y_{h}^{m}\|_{L^{2}}).$
[…]
$\interleave y-y_{h}^{m}\interleave_{\mathrm{DG}}\leq C(\interleave y-P_{h}y\interleave_{\mathrm{DG}}+\|u-u_{h}\|_{L^{2}}).$
Combine the above inequalities, we have
$\|u-u_{h}\|_{L^{2}}+\interleave y-y_{h}^{m}\interleave_{\mathrm{DG}}+\interleave p-p_{h}^{m}\interleave_{\mathrm{DG}}$
$\leq$ $C(\|(I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p)\|_{L^{2}}+\|u-\Pi_{h}^{u}u\|_{L^{2}}+\interleave p-P_{h}p\interleave_{\mathrm{DG}}+\interleave y-P_{h}y\interleave_{\mathrm{DG}}).$
Notice that if we substitute all the $\interleave\cdot\interleave_{\mathrm{DG}}$ by $\|\cdot\|_{L^{2}}$, this inequality still holds.
For Lipschitz continuous function $v\in U,$ $\|v-\Pi_{h}^{u}v\|\leq Ch_{u}$. Moreover, by Theorem (1), we have
$\interleave p-P_{h}p\interleave_{\mathrm{DG}}+\interleave y-P_{h}y\interleave_{\mathrm{DG}}\leq Ch^{m},$
and also
$\|p-P_{h}p\|_{L^{2}}+\|y-P_{h}y\|_{L^{2}}\leq Ch^{m+1}.$
Then the proof is completed. ∎
3.2. Super-convergence Analysis.
We mention here that the order of $h_{u}$ on the right-hand side of Theorem 3 can be improved to 3/2{3}/{2} if we concentrate on the case where
$U_{ad}=\{v\in L^{2}({\Omega}_{u})\mid v\geq\gamma\ \mathrm{a.e.\ in}\ {\Omega}_{u}\},$
where $\gamma$ is a constant. To derive the higher order convergence, we need more assumptions on regularity as follows. Consider two disjoint subsets of $\Omega_{u}$.
${\Omega}_{u}^{+}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u^{*}|_{K_{u}}>\gamma\},\quad{\Omega}_{u}^{0}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u^{*}|_{K_{u}}=\gamma\}.$
Define $\Omega_{u}^{b}:=\Omega_{u}\setminus(\Omega_{u}^{+}\cup\Omega_{u}^{0}).$
Assumption 4.
The measure of $\Omega_{u}^{b}$ such that $m(\Omega_{u}^{b})\leq Ch_{u},$ where $C>0$ is a constant.
Assumption 4 can be easily achieved for the case that the exact solution $u^{*}$ has the form $u^{*}=\max\{w,\gamma\}$ with the function $w$ smooth enough. For sufficiently smooth $w$, the set $\{w=\gamma\}$ would be a smooth rectifiable curve, which results in $m(\Omega_{u}^{b})\leq Ch_{u}$ by the quasi-uniformity of $\mathcal{T}_{h}^{u}.$
Consider the space $H^{1}(\mathcal{T}_{h}^{u})$ equipped with the energy norm $\|\cdot\|_{\mathrm{DG}}^{u}$, which has the same expression as the norm $\|\cdot\|_{\mathrm{DG}}$ on $H^{1}(\mathcal{T}_{h})$. We want to add more regularity to the operator $B^{*}:L^{2}(\Omega)\rightarrow L^{2}(\Omega_{u}).$
Assumption 5.
The operator $B^{*}$ satisfies $B^{*}V_{h}^{m}\subset H^{1}(\mathcal{T}_{h}^{u})$ and $B^{*}|_{V_{h}^{m}}:V_{h}^{m}\rightarrow H^{1}(\mathcal{T}_{h}^{u})$ is bounded.
Now we use these two assumptions to prove the super-close property of $u_{h}^{*}$ and $\Pi_{h}^{u}u^{*}$.
Lemma 3.
[…]
$\|u_{h}-\Pi_{h}^{u}u\|_{L^{2}}\leq C(h_{u}^{\frac{3}{2}}+h^{m+1}).$
Proof.
Consider the intermediate state as follows.
(3.5) $a_{h}(y_{\Pi}^{m},w_{h}^{m})=l_{h}[f+B\Pi_{h}^{u}u,\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$a_{h}(q_{h}^{m},p_{\Pi}^{m})=l_{h}[g^{\prime}(y_{\Pi}^{m}),0](q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m}.$
$y_{\Pi}^{m},p_{\Pi}^{m}\in V_{h}^{m}.$ From (2.12) and (3.5), we have
(3.6) $a_{h}(y_{\Pi}^{m}-y_{h}^{m},w_{h}^{m})=(B(\Pi_{h}^{u}u-u_{h}),w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
$a_{h}(q_{h}^{m},p_{\Pi}^{m}-p_{h}^{m})=(g^{\prime}(y_{\Pi}^{m})-g^{\prime}(y_{h}^{m}),q_{h}^{m}),\quad\forall q_{h}^{m})\in V_{h}^{m}.$
Take $w_{h}^{m}=p_{\Pi}^{m}-p_{h}^{m},q_{h}^{m}=y_{\Pi}^{m}-y_{h}^{m}$ in (3.6) and subtract the resulting equations to find
$(B^{*}(p_{\Pi}^{m}-p_{h}^{m}),\Pi_{h}^{u}u-u_{h})=(g^{\prime}(y_{\Pi}^{m})-g^{\prime}(y_{h}^{m}),y_{\Pi}^{m}-y_{h}^{m})\geq 0.$
Therefore, we have
$\alpha\|u_{h}-\Pi_{h}^{u}u\|_{L^{2}}^{2}\leq$ $(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-\Pi_{h}^{u}u)-(j^{\prime}(\Pi_{h}^{u}u)+B^{*}p_{\Pi}^{m},u_{h}-\Pi_{h}^{u}u)$
$=$ $(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-\Pi_{h}^{u}u)+(j^{\prime}(u)+B^{*}p,\Pi_{h}^{u}u-u_{h})$
$+(j^{\prime}(\Pi_{h}^{u}u)-j^{\prime}(u),\Pi_{h}^{u}u-u_{h})+(B^{*}p_{\Pi}^{m}-B^{*}p,\Pi_{h}^{u}u-u_{h})$
$\leq$ $0+(j^{\prime}(u)+B^{*}p,u-u_{h})+(j^{\prime}(u)+B^{*}p,\Pi_{h}^{u}u-u)$
$+(j^{\prime}(\Pi_{h}^{u}u)-j^{\prime}(u),\Pi_{h}^{u}u-u_{h})+(B^{*}p_{\Pi}^{m}-B^{*}p,\Pi_{h}^{u}u-u_{h})$
$\leq$ $0+0+I_{1}+I_{2}+I_{3}.$
Here $I_{1},I_{2},I_{3}$ are the last three terms respectively.
By the results in [24](Lemma 4.3.1), we know that under Assumption 4,
$I_{1}\leq Ch_{u}^{3},\quad I_{2}\leq Ch_{u}^{2}\|\Pi_{h}^{u}u-u_{h}\|_{L^{2}},$
where $C$ depends on $\|u\|_{W^{1,\infty}}$ and $\|j^{\prime}(u)+B^{*}p\|_{W^{1,\infty}}.$ Next we will give upper bounds of $\|y_{\Pi}^{m}-y\|_{L^{2}}$ and $\|p_{\Pi}^{m}-p\|_{L^{2}}$ to estimate $I_{3}$.
[…]
$\|\Pi_{h}^{u}u-u\|_{-\mathrm{DG}}^{u}$ $=\sup_{v\in H^{1}(\mathcal{T}_{h}^{u})}\frac{(\Pi_{h}^{u}u-u,v)}{\|v\|_{\mathrm{DG}}^{u}}=\sup_{v\in H^{1}(\mathcal{T}_{h}^{u})}\frac{(\Pi_{h}^{u}u-u,v-\Pi_{h}^{u}v)}{\|v\|_{\mathrm{DG}}^{u}}$
$\;\leq\sup_{v\in H^{1}(\mathcal{T}_{h}^{u})}\frac{C\sum_{\tau_{u}\in\mathcal{T}_{h}^{u}}h_{\tau_{u}}^{2}|u|_{1,\tau_{u}}|v|_{1,\tau_{u}}}{\|v\|_{\mathrm{DG}}^{u}}\leq Ch_{u}^{2}\|u\|_{H^{1}}.$
Recall the definitions of $P_{h}y,P_{h}p$ in Theorem 3, we have
$a_{h}(y_{\Pi}^{m}-P_{h}y,y_{\Pi}^{m}-P_{h}y)=(B(\Pi_{h}^{u}u-u),y_{\Pi}^{m}-P_{h}y),$
$a_{h}(p_{\Pi}^{m}-P_{h}p,p_{\Pi}^{m}-P_{h}p)=(g^{\prime}(y_{\Pi}^{m})-g^{\prime}(y),p_{\Pi}^{m}-P_{h}p).$
By the stability of $a_{h}$ on $(V_{h}^{m},\|\cdot\|_{\mathrm{DG}})$,
$c\|y_{\Pi}^{m}-P_{h}y\|_{\mathrm{DG}}^{2}\leq a_{h}(y_{\Pi}^{m}-P_{h}y,y_{\Pi}^{m}-P_{h}y)\leq\|\Pi_{h}^{u}u-u\|_{-\mathrm{DG}}^{u}\|B^{*}(y_{\Pi}^{m}-P_{h}y)\|_{\mathrm{DG}}^{u}.$
Using Assumption 5, we have
$\|y_{\Pi}^{m}-P_{h}y\|_{\mathrm{DG}}\leq C\|\Pi_{h}^{u}u-u\|_{-\mathrm{DG}}^{u}\leq Ch_{u}^{2}\|u\|_{H^{1}}.$
By (2.8), we know that
$\|y_{\Pi}^{m}-y\|_{L^{2}}\leq C\|y_{\Pi}^{m}-P_{h}y\|_{\mathrm{DG}}+\|P_{h}y-y\|_{L^{2}}\leq C(h_{u}^{2}+h^{m+1}).$
Similarly, we have
$c\|p_{\Pi}^{m}-P_{h}p\|_{L^{2}}^{2}\leq a_{h}(p_{\Pi}^{m}-P_{h}p,p_{\Pi}^{m}-P_{h}p)\leq\|g^{\prime}(y_{\Pi}^{m})-g^{\prime}(y)\|_{L^{2}}\|p_{\Pi}^{m}-P_{h}p\|_{L^{2}},$
then
$\|p_{\Pi}^{m}-P_{h}p\|_{L^{2}}\leq C\|y_{\Pi}^{m}-y\|_{L^{2}}\leq C(h_{u}^{2}+h^{m+1}).$
Finally
$\|p_{\Pi}^{m}-p\|_{L^{2}}\leq\|p_{\Pi}^{m}-P_{h}p\|_{L^{2}}+\|P_{h}p-p\|_{L^{2}}\leq C(h_{u}^{2}+h^{m+1}).$
We now have
$I_{1}+I_{2}+I_{3}\leq Ch_{u}^{3}+C(h_{u}^{2}+h^{m+1})\|\Pi_{h}^{u}u-u_{h}\|_{L^{2}}.$
Therefore $\|\Pi_{h}^{u}u-u_{h}\|_{L^{2}}\leq C(h_{u}^{\frac{3}{2}}+h^{m+1}).$
Remark 3.
For the case
$U_{ad}=\{v\in L^{2}({\Omega}_{u})\mid\gamma\leq v\leq\Gamma\ \mathrm{a.e.\ in}\ {\Omega}_{u}\},$
where $\gamma<\Gamma$ are constants, take
$\Omega_{u}^{+}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid\gamma<u_{h}^{*}|_{K_{u}}<\Gamma\},\quad\Omega_{u}^{0}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u_{h}^{*}|_{K_{u}}=\gamma\mathrm{\ or\ }\Gamma\}.$
The result in Lemma 3 still holds.
In order to make use of Lemma 3, we need to find a recovery operator $R_{h}:U\rightarrow U$ to get a sharp estimate of $\|u^{*}-R_{h}u^{*}_{h}\|_{L^{2}}.$ More precisely, we have the inequality
$\|u^{*}-R_{h}u^{*}_{h}\|_{L^{2}}\leq\|u^{*}-R_{h}(\Pi_{h}^{u}u^{*})\|_{L^{2}}+\|R_{h}\|\|\Pi_{h}^{u}u^{*}-u_{h}^{*}\|_{L^{2}}.$
Given the estimate for $\|\Pi_{h}^{u}u^{*}-u_{h}^{*}\|_{L^{2}}$, it remains to select a bounded linear operator $R_{h}$ ensuring that $\|u^{*}-R_{h}(\Pi_{h}^{u}u^{*})\|_{L^{2}}$ is of a higher order.
In particular, when $\Omega_{u}\subset\mathbb{R}^{2}$, the recovery $R_{h}v(z)$ for the node $z=(x_{0},y_{0})\in\Omega_{u}$ can be calculated by a kind of Z-Z patch recovery. Let $K^{u}_{1},K^{u}_{2},\cdots,K^{u}_{m}$ be the elements in $\mathcal{T}_{h}^{u}$ containing $z$. Suppose the three nodes of $K^{u}_{i}$ are $(x_{0},y_{0}),(x_{i},y_{i})$ and $(x_{i+1},y_{i+1}),i=1,2,\ldots,m$. Define $R_{h}v(z):=a+bx_{0}+cy_{0}$, where $(a,b,c)$ is the solution to the following linear system.
[…]
Here ($\overline{x_{i}},\overline{y_{i}})$ is the barycenter of $K^{u}_{i}$. $|k_{i}|=m(K^{u}_{i}),v_{i}=\int_{K^{u}_{i}}v\mathrm{d}x.$
Consequently, we can define the operator $R_{h}:U\rightarrow\mathbb{P}_{1}(\mathcal{T}_{h}^{u})$ (the piecewise linear continuous function space on $\mathcal{T}_{h}^{u}$) without confusion. From [3], we know that $R_{h}$ is bounded and if $u^{*}\in W^{1,\infty}(\Omega_{u}),u^{*}|_{\Omega_{u}^{+}}\in H^{2}(\Omega_{u}^{+}),$ then there exists a constant $C=C(u^{*})$ such that
$\|u^{*}-R_{h}(\Pi_{h}^{u}u^{*})\|_{L^{2}}=\|u^{*}-R_{h}u^{*}\|_{L^{2}}\leq Ch_{u}^{2}.$
[…]
Theorem 4.
[…]
$\|u-R_{h}u_{h}\|_{L^{2}}\leq C(h_{u}^{\frac{3}{2}}+h^{m+1}).$
4. A Posteriori Estimates
4.1. A General Result
This section conducts an a posteriori error analysis for the problem (2.12). Analogous to the proof of Theorem 3, we will also construct two intermediate states. This recasts the error estimate for (2.12) into an analysis of the DG method’s error for the underlying elliptic equation, thereby permitting the application of the established results from Theorem 2.
Theorem 5.
[…]
$\|u-u_{h}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}+\|p-p_{h}^{m}\|_{L^{2}}$
$\leq$ $C(\eta_{0}+\eta_{1}(y_{h}^{m})+\eta_{2}(y_{h}^{m})+\eta_{3}(y_{h}^{m})+\eta_{1}(p_{h}^{m})+\eta_{2}(p_{h}^{m})+\eta_{3}(p_{h}^{m})),$
where
$\eta_{0}^{2}=\|(I-\Pi_{h}^{u})(j^{\prime}(u_{h})+B^{*}p_{h}^{m})\|_{L^{2}}^{2},$
$\eta_{1}^{2}(y_{h}^{m})=\sum_{K\in\mathcal{T}^{h}}\widetilde{h_{K}}^{4}\|f+Bu_{h}+\nabla\cdot(A\nabla y_{h}^{m})\|_{L^{2}(K)}^{2},$
$\eta_{1}^{2}(p_{h}^{m})=\sum_{K\in\mathcal{T}^{h}}\widetilde{h_{K}}^{4}\|g^{\prime}(y_{h}^{m})+\nabla\cdot(A\nabla p_{h}^{m})\|_{L^{2}(K)}^{2},$
$\eta_{2},\eta_{3}$ are defined by (2.10).
Proof.
Construct the intermediate states as follows.
$-\nabla\cdot(A\nabla Sy)=f+Bu_{h}\quad Sy|_{\partial\Omega}=\phi,$
$-\nabla\cdot(A\nabla Sp)=g^{\prime}(y_{h}^{m}),\quad Sp|_{\partial\Omega}=0,$
which means that
(4.1) $a_{h}(Sy,w_{h})=l_{h}[f+Bu_{h},\phi](w_{h}),\quad\forall w_{h}\in V_{h},$
$a_{h}(q_{h},Sp)=l_{h}[g^{\prime}(y_{h}^{m}),0](q_{h}),\quad\forall q_{h}\in V_{h}.$
From (2.11) and (4.1), we have
(4.2) $a_{h}(y-Sy,w_{h})=(B(u-u_{h}),w_{h}),\quad\forall w_{h}\in V_{h},$
$a_{h}(q_{h},p-Sp)=(g^{\prime}(y)-g^{\prime}(y_{h}^{m}),q_{h}),\quad\forall q_{h}\in V_{h}.$
Take $w_{h}=p-Sp,q_{h}=y-Sy$ in (4.2) and subtract the resulting equations to find
$(B^{*}(p-Sp),u-u_{h})=(g^{\prime}(y)-g^{\prime}(y_{h}^{m}),y-Sy).$
This implies
(4.3) $(B^{*}(p-Sp),u-u_{h})\geq\beta\|y-Sy\|_{L^{2}}+(g^{\prime}(Sy)-g^{\prime}(y_{h}^{m}),y-Sy).$
In addition, by (3.4) in the proof of Theorem 3, we know that
(4.4) $(B^{*}(p_{h}^{m}-p),u-u_{h})\geq\alpha\|u-u_{h}\|_{L^{2}}^{2}+(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-v_{h}),\quad\forall v_{h}\in U_{adh}.$
Take $v_{h}=\Pi_{h}^{u}u.$
Combine (4.3) and (4.4), we have
$\alpha\|u-u_{h}\|_{L^{2}}^{2}+\beta\|y-Sy\|_{L^{2}}^{2}$
$\leq$ $-(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u-\Pi_{h}^{u}u)-(B^{*}(Sp-p_{h}^{m}),u-u_{h})+(g^{\prime}(y_{h}^{m})-g^{\prime}(Sy),y-Sy)$
$=$ $-((I-\Pi_{h}^{u})(j^{\prime}(u_{h})+B^{*}p_{h}^{m}),u-\Pi_{h}^{u}u)$
$-(B^{*}(Sp-p_{h}^{m}),u-u_{h})+(g^{\prime}(y_{h}^{m})-g^{\prime}(Sy),y-Sy)$
$=$ $-((I-\Pi_{h}^{u})(j^{\prime}(u_{h})+B^{*}p_{h}^{m}),u-u_{h})$
$-(B^{*}(Sp-p_{h}^{m}),u-u_{h})+(g^{\prime}(y_{h}^{m})-g^{\prime}(Sy),y-Sy)$
$\leq$ $C(\eta_{0}^{2}+\|B^{*}(Sp-p_{h}^{m})\|_{L^{2}}^{2}+\|Sy-y_{h}^{m}\|_{L^{2}}^{2})+\frac{\alpha}{2}\|u-u_{h}\|_{L^{2}}^{2}+\frac{\beta}{2}\|y-Sy\|_{L^{2}}^{2}.\quad\quad\quad\,$
So we have
$\|u-u_{h}\|_{L^{2}}+\|y-Sy\|_{L^{2}}\leq C(\eta_{0}+\|Sp-p_{h}^{m}\|_{L^{2}}+\|Sy-y_{h}^{m}\|_{L^{2}}).$
[…]
$\|y-Sy\|_{L^{2}}\leq C\|u-u_{h}\|_{L^{2}},\quad\|p-Sp\|_{L^{2}}\leq C\|y-y_{h}^{m}\|_{L^{2}}.$
Thus
$\|p-p_{h}^{m}\|_{L^{2}}\leq C(\|Sp-p_{h}^{m}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}),$
$\|y-y_{h}^{m}\|_{L^{2}}\leq C(\|Sy-y_{h}^{m}\|_{L^{2}}+\|u-u_{h}\|_{L^{2}}).$
Combine the three inequalities above, we have
$\|u-u_{h}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}+\|p-p_{h}^{m}\|_{L^{2}}\leq C(\eta_{0}+\|Sp-p_{h}^{m}\|_{L^{2}}+\|Sy-y_{h}^{m}\|_{L^{2}}).$
Moreover, by Theorem (2), we have
$\|Sp-p_{h}^{m}\|_{L^{2}}\leq C(\eta_{1}(p_{h}^{m})+\eta_{2}(p_{h}^{m})+\eta_{3}(p_{h}^{m})),$
$\|Sy-y_{h}^{m}\|_{L^{2}}\leq C(\eta_{1}(y_{h}^{m})+\eta_{2}(y_{h}^{m})+\eta_{3}(y_{h}^{m})),$
which leads to the conclusion. ∎
4.2. Sharp A Posteriori Estimate.
We concentrate on the case where
$U_{ad}=\{v\in L^{2}({\Omega}_{u})\mid v\geq\gamma\ \mathrm{a.e.\ in}\ {\Omega}_{u}\},$
where $\gamma$ is a constant. We divide ${\Omega}_{u}$ into two disjoint subsets.
$\widehat{\Omega}_{u}^{+}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u_{h}^{*}|_{K_{u}}>\gamma\},\quad\widehat{\Omega}_{u}^{0}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u_{h}^{*}|_{K_{u}}=\gamma\}.$
Consider a subset of $\widehat{\Omega}_{u}^{0}$
$\widehat{\Omega}_{u1}^{0}=\{x\in\widehat{\Omega}_{u}^{0}\mid j^{\prime}(u_{h}^{*})+B^{*}p_{h}^{m*}\geq 0\}.$
Theorem 6.
[…]
$\eta_{*}+\|u-u_{h}\|_{L^{2}}+\|y-y_{h}^{m}\|_{L^{2}}+\|p-p_{h}^{m}\|_{L^{2}}$
$\leq$ $C(\widetilde{\eta_{0}}+\eta_{1}(y_{h}^{m})+\eta_{2}(y_{h}^{m})+\eta_{3}(y_{h}^{m})+\eta_{1}(p_{h}^{m})+\eta_{2}(p_{h}^{m})+\eta_{3}(p_{h}^{m})),$
where
$\eta_{*}\,=\|(I-\Pi_{h}^{u})(j^{\prime}(u)+B^{*}p)\|_{L^{2}(\widehat{\Omega}_{u}^{+})},\quad\widetilde{\eta_{0}}\,=\|j^{\prime}(u_{h})+B^{*}p_{h}^{m}\|_{L^{2}((\widehat{\Omega}_{u1}^{0})^{c})}.$
$\eta_{1},\eta_{2},\eta_{3}$ are defined in Theorem 5. Here $(\widehat{\Omega}_{u1}^{0})^{c}=\Omega_{u}\setminus\widehat{\Omega}_{u1}^{0}.$
Proof.
For $\forall K\in\mathcal{T}_{h}^{u},$ take $v_{h}=\gamma\chi_{K}+u_{h}\chi_{{\Omega}_{u}\setminus K}\in U_{adh}.$ Here $\chi_{K}$ is the indicator function of $K$. We know by (2.12) that
$(\Pi_{h}^{u}(j^{\prime}(u_{h})+B^{*}p_{h}^{m}),v_{h}-u_{h})=(j^{\prime}(u_{h})+B^{*}p_{h}^{m},v_{h}-u_{h})\geq 0,$
which implies $\Pi_{h}^{u}(j^{\prime}(u_{h})+B^{*}p_{h}^{m})=0$ on $\widehat{\Omega}_{u}^{+}.$
$\eta_{*}=$ $\|j^{\prime}(u)+B^{*}p-\Pi_{h}^{u}(j^{\prime}(u)+B^{*}p)\|_{L^{2}(\widehat{\Omega}_{u}^{+})}$
$\leq$ $\|j^{\prime}(u)+B^{*}p-j^{\prime}(u_{h})-B^{*}p_{h}^{m}\|_{L^{2}(\widehat{\Omega}_{u}^{+})}+\|(I-\Pi_{h}^{u})(j^{\prime}(u_{h})+B^{*}p_{h}^{m})\|_{L^{2}(\widehat{\Omega}_{u}^{+})}$
$+\|\Pi_{h}^{u}(j^{\prime}(u_{h})+B^{*}p_{h}^{m}-j^{\prime}(u)-B^{*}p)\|_{L^{2}(\widehat{\Omega}_{u}^{+})}$
$\leq$ $C(\|j^{\prime}(u)+B^{*}p-j^{\prime}(u_{h})-B^{*}p_{h}^{m}\|_{L^{2}(\widehat{\Omega}_{u}^{+})}+\|j^{\prime}(u_{h})+B^{*}p_{h}^{m}\|_{L^{2}(\widehat{\Omega}_{u}^{+})})$
$\leq$ $C(\|u-u_{h}\|_{L^{2}(\widehat{\Omega}_{u}^{+})}+\|p-p_{h}^{m}\|_{L^{2}(\widehat{\Omega}_{u}^{+})}+\|j^{\prime}(u_{h})+B^{*}p_{h}^{m}\|_{L^{2}(\widehat{\Omega}_{u}^{+})})$
[…]
Notice that $j^{\prime}(u_{h})+B^{*}p_{h}^{m}\geq 0$ on $\widehat{\Omega}_{u1}^{0}$, we have
$(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-u)$
$=$ $(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-u)_{L^{2}(\widehat{\Omega}_{u1}^{0})}+(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-u)_{L^{2}((\widehat{\Omega}_{u1}^{0})^{c})}$
$\leq$ $0+C\widetilde{\eta_{0}}^{2}+\frac{\alpha}{2}\|u_{h}-u\|_{L^{2}((\widehat{\Omega}_{u1}^{0})^{c})}^{2}\leq C\widetilde{\eta_{0}}^{2}+\frac{\alpha}{2}\|u_{h}-u\|_{L^{2}}^{2}.$
Follow the proof of Theorem 5 but taking $v_{h}=u_{h}$ in (4.4),
$\alpha\|u-u_{h}\|_{L^{2}}^{2}+\beta\|y-Sy\|_{L^{2}}^{2}$
$\leq$ $(j^{\prime}(u_{h})+B^{*}p_{h}^{m},u_{h}-u)-(B^{*}(Sp-p_{h}^{m}),u-u_{h})+(g^{\prime}(y_{h}^{m})-g^{\prime}(Sy),y-Sy)$
$\leq$ $C(\widetilde{\eta_{0}}^{2}+\|Sp-p_{h}^{m}\|_{L^{2}}^{2}+\|Sy-y_{h}^{m}\|_{L^{2}}^{2})+\frac{3\alpha}{4}\|u-u_{h}\|_{L^{2}}^{2}+\frac{\beta}{2}\|y-Sy\|_{L^{2}}^{2}.$
So we have
$\|u-u_{h}\|_{L^{2}}+\|y-Sy\|_{L^{2}}\leq C(\widetilde{\eta_{0}}+\|Sp-p_{h}^{m}\|_{L^{2}}+\|Sy-y_{h}^{m}\|_{L^{2}}).$
Using the estimate of $\eta_{*}$ as above, we can finish the proof by the same procedure in Theorem 5. ∎
Remark 4.
For the case where
$U_{ad}=\{v\in L^{2}({\Omega}_{u})\mid\gamma\leq v\leq\Gamma\ \mathrm{a.e.\ in}\ {\Omega}_{u}\},$
where $\gamma<\Gamma$ are constants, take
$\widehat{\Omega}_{u}^{+}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid\gamma<u_{h}^{*}|_{K_{u}}<\Gamma\},\quad\widehat{\Omega}_{u}^{0}=\{K_{u}\in\mathcal{T}_{h}^{u}\mid u_{h}^{*}|_{K_{u}}=\gamma\mathrm{\ or\ }\Gamma\},$
and
[…]
$\{x\in\widehat{\Omega}_{u}^{0}\mid u_{h}^{*}(x)=\Gamma,j^{\prime}(u_{h}^{*})+B^{*}p_{h}^{m*}\leq 0\}.$
The inequality in Theorem 6 still holds.
5. Numerical Results
In thus section, we provide some numerical examples to validate the derivation in the previous sections. In Section 5.1, we verify the convergence order of a priori estimate proved in Theorem 3 and Theorem 4. In Section 5.2, we study the case that the control variable has sufficiently smooth data. We build another approximation to the optimality condition (2.11) without using $\mathcal{T}_{h}^{u}$, which means that we do not discretize the control variable. In Section 5.3, we demonstrate the efficiency of a posteriori estimate proved in Theorem 5.
5.1. Examples for Convergence Order.
[…]
Example 1. Take $\Omega_{u}=\Omega=(0,1)^{2},B=\mathrm{Id}.$
$\min_{u\in L^{2}(\Omega_{u})}\frac{1}{2}\|y-y_{d}\|_{L^{2}}^{2}+\frac{1}{2}\|u-u_{d}\|_{L^{2}}^{2},$
$\mathrm{s.t.}$ $-\Delta y=f+u,\quad y|_{\partial\Omega}=y_{d}|_{\partial\Omega},$
$u\geq 0,\quad\mathrm{a.e.\ in}\ \Omega_{u}.$
where
$y_{d}=0,\quad u_{d}=1-\mathrm{sin}\frac{\pi x_{1}}{2}-\mathrm{sin}\frac{\pi x_{2}}{2},$
$p=\mathrm{sin}(\pi x_{1})\mathrm{sin}(\pi x_{2}),\quad f=4\pi^{4}p-\mathrm{max}\{u_{d}-p,0\}.$
The exact solution
$y^{*}=2\pi^{2}p+y_{d},\quad u^{*}=\mathrm{max}\{u_{d}-p,0\},\quad p^{*}=p.$
Figure 1. Error of Example 1 for $m=1.$ Take $h_{u}=h$(left) and $h_{u}=4h^{2}$(right).
Figure 2. Error of Example 1 for $m=2.$ Take $h_{u}=4h^{2}$(left) and $h_{u}=16h^{3}$(right).
Figure 3. Error of Example 1 for $m=3$. Take $h_{u}=16h^{3}$.
Figure 4. Recovery error of Example 1. Take $h_{u}=h.$
Example 2. Take $\Omega_{u}=\Omega=(0,1)^{2},B=\displaystyle\frac{1}{2}\mathrm{Id}.$
$\min_{u\in L^{2}(\Omega_{u})}\frac{1}{2}\|y-y_{d}\|_{L^{2}}^{2}+\frac{1}{2}\|u-u_{d}\|_{L^{2}}^{2}+\frac{4}{3}\|u-u_{d}\|_{L^{3}}^{3},$
$\mathrm{s.t.}$ $-\Delta y=f+\frac{1}{2}u,\quad y|_{\partial\Omega}=y_{d}|_{\partial\Omega},$
$0\leq u\leq 1,\quad\mathrm{a.e.\ in}\ \Omega_{u}.$
where
$u_{d}=\frac{3}{2}-x_{1}-x_{2},\quad p=\mathrm{sin}(\pi x_{1})\mathrm{sin}(\pi x_{2}),$
$y_{d}=0,\quad f=4\pi^{4}p-\frac{1}{2}Pr_{[0,1]}(u_{d}-s(p)).$
The exact solution
$y^{*}=2\pi^{2}p+y_{d},\quad u^{*}=Pr_{[0,1]}(u_{d}-s(p)),\quad p^{*}=p,$
where $s(p)$ is an implicit function satisfies
$s(p)+4s(p)^{2}\mathrm{sgn}(s(p))=\frac{1}{2}p,$
and $Pr_{[0,1]}(s):=\min\{\max\{s,0\},1\}.$
Figure 5. Error of Example 2 for $m=1.$ Take $h_{u}=h$(left) and $h_{u}=4h^{2}$(right).
Figure 6. Error of Example 2 for $m=2$. Take $h_{u}=4h^{2}$(left) and $h_{u}=16h^{3}$(right).
Figure 7. Error of Example 2 for $m=3$. Take $h_{u}=16h^{3}.$
Figure 8. Recovery error of Example 2. Take $h_{u}=h.$
5.2. Smooth Example.
Notice that in Theorem 3 and Theorem 4, the order of $h_{u}$ is higher compared to $h$, which leads to more refinement on $\mathcal{T}_{h}^{u}$. In order to optimize computational efficiency and minimize memory cost, we can calculate $u$ without discretization and eliminate $\mathcal{T}_{h}^{u}$, i.e. we consider an approximation to (2.11) better than (2.12) as follows.
(5.1) $\left\{\begin{aligned} &a_{h}(\widetilde{y}_{h}^{m\ast},w_{h}^{m})=l_{h}[f+B\widetilde{u}^{*},\phi](w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},\\ &a_{h}(q_{h}^{m},\widetilde{p}_{h}^{m\ast})=l_{h}[g^{\prime}(\widetilde{y}_{h}^{m\ast}),0](q_{h}^{m}),\quad\forall q_{h}^{m}\in V_{h}^{m},\\ &(j^{\prime}(\widetilde{u}^{*})+B^{*}\widetilde{p}_{h}^{m*},v-\widetilde{u}^{*})\geq 0,\quad\forall v\in U_{ad}.\end{aligned}\right$.
We denote the solution to problem (5.1) by ($\widetilde{y}_{h}^{m\ast},\widetilde{u}^{*},\widetilde{p}_{h}^{m*}$). Follow the proof of Theorem 3, we can take $v_{h}=u^{*}$ to derive
$\|u^{*}-\widetilde{u}^{*}\|_{L^{2}}^{2}+\|\widetilde{y}_{h}^{m*}-P_{h}{y}\|_{L^{2}}^{2}\leq C(\|p^{*}-P_{h}p\|_{L^{2}}^{2}+\|y^{*}-P_{h}y\|_{L^{2}}^{2})$
where $C$ depends on $\alpha,\beta.$ Consequently, we have
$\|u^{*}-\widetilde{u}^{*}\|_{L^{2}}+\interleave y^{*}-\widetilde{y}_{h}^{m*}\interleave_{\mathrm{DG}}+\interleave p^{*}-\widetilde{p}_{h}^{m*}\interleave_{\mathrm{DG}}\leq Ch^{m},$
and
$\|u^{*}-\widetilde{u}^{*}\|_{L^{2}}+\|y^{*}-\widetilde{y}_{h}^{m*}\|_{L^{2}}+\|p^{*}-\widetilde{p}_{h}^{m*}\|_{L^{2}}\leq Ch^{m+1}.$
In each iteration, we rewrite the third equation in (2.14) as
$\widetilde{u}_{\mkern 1.0mun\mkern-2.0mu+\mkern-2.0mu1}=Pr_{U_{ad}}(\widetilde{u}_{n}-\rho_{n}(j^{\prime}(\widetilde{u}_{n})+B^{*}\widetilde{p}_{hn}^{m})).$
We can prove that the linear convergence rate still holds. The smoothness of $u^{*}$ permits high-order accurate numerical integration of $\widetilde{u}_{n}$-dependent integrands by evaluating directly at the quadrature points of $\mathcal{T}_{h}.$ Consequently, we only need to store the values of $\widetilde{u}_{n}$ at these quadrature points. This modification serves to significantly improve computational efficiency and offers the advantage of reduced memory consumption.
Example 3. Take $\Omega_{u}=\Omega=(0,1)^{2},B=\mathrm{Id}.$
$\min_{u\in L^{2}(\Omega_{u})}\frac{1}{2}\|y-y_{d}\|_{L^{2}}^{2}+\frac{1}{2}\|u-u_{d}\|_{L^{2}}^{2},$
$\mathrm{s.t.}$ $-\Delta y=f+u,\quad y|_{\partial\Omega}=y_{d}|_{\partial\Omega},\quad\int_{\Omega_{u}}u\mathrm{d}x\geq 0.$
where
$u_{d}=1-2x_{1}-2x_{2},\quad p=\mathrm{sin}(\pi x_{1})\mathrm{sin}(\pi x_{2}),$
$y_{d}=0,\quad f=4\pi^{4}p-u_{d}+p+\min\{0,\overline{u_{d}-p}\}.$
The exact solution
$y^{*}=2\pi^{2}p+y_{d},\quad u^{*}=u_{d}-p-\min\{0,\overline{u_{d}-p}\},\quad p^{*}=p,$
here $\overline{u_{d}-p}=\displaystyle\frac{1}{m(\Omega)}\int_{\Omega}(u_{d}-p)\mathrm{d}x.$
Figure 9. Error of Example 3 for $m=1.$
Figure 10. Error of Example 3 for $m=2.$
Figure 11. Error of Example 3 for $m=3.$
5.3. A Posteriori Estimator.
We will show the efficiency of a posteriori error estimator provided in Theorem 5. We will plot the ratio of a posteriori error indicator and local error for three variables $(y,u,p)$. For the control variable $u$, a posteriori error estimators
$e_{K_{u}}^{2}:=\frac{\|j^{\prime}(u^{*}_{h})+B^{*}p_{h}^{m*}\|^{2}_{L^{2}(K_{u})}}{\|u^{*}-u_{h}^{*}\|^{2}_{L^{2}(K_{u})}},\quad\forall K_{u}\in\mathcal{T}_{h}^{u}.$
For the variables $y,p$ and $K\in\mathcal{T}_{h}$, we write
$\eta_{K_{y}}^{2}={\widetilde{h_{K}}^{4}\|f+Bu_{h}^{*}+\nabla\cdot(A\nabla y_{h}^{m*})\|_{L^{2}(K)}^{2}+\sum_{e\in\mathcal{E}_{K}\cap\mathcal{E}_{h}}\widetilde{h_{e}}\|[y_{h}^{m*}]\|_{L^{2}(e)}^{2}+\sum_{e\in\mathcal{E}_{K}\cap\mathcal{E}_{h}^{I}}\widetilde{h_{e}}^{3}\|[A\nabla y_{h}^{m*}]\|_{L^{2}(e)}^{2}},$
$\eta_{K_{p}}^{2}={\widetilde{h_{K}}^{4}\|g^{\prime}(y_{h}^{m*})+\nabla\cdot(A\nabla p_{h}^{m*})\|_{L^{2}(K)}^{2}+\sum_{e\in\mathcal{E}_{K}\cap\mathcal{E}_{h}}\widetilde{h_{e}}\|[p_{h}^{m*}]\|_{L^{2}(e)}^{2}+\sum_{e\in\mathcal{E}_{K}\cap\mathcal{E}_{h}^{I}}\widetilde{h_{e}}^{3}\|[A\nabla p_{h}^{m*}]\|_{L^{2}(e)}^{2}}.$
We define
$e_{K_{y}}^{2}:=\frac{\eta_{K_{y}}^{2}}{\|y^{*}-y_{h}^{m*}\|_{L^{2}(K)}^{2}},\quad e_{K_{p}}^{2}:=\frac{\eta_{K_{p}}^{2}}{\|p^{*}-p_{h}^{m*}\|_{L^{2}(K)}^{2}},\quad\forall K\in\mathcal{T}_{h}.$
In addition, these expressions are defined as 0 whenever their denominators are 0.
Example 4. We will present some results of $e_{K_{y}},e_{K_{p}}$ and $e_{K_{u}}$ for the problem in Example 1. We can see that the maximum and minimum values of $e_{K_{y}}$ and $e_{K_{p}}$ show no significant variation with the refinement of $\mathcal{T}_{h}$, illustrating the efficiency $\eta_{K_{y}}$ and $\eta_{K_{p}}$.
$h$ 1/4 1/8 1/16 1/32 1/64 1/128
$\max_{K}e_{K_{y}}$ 9.656 18.10 20.88 24.28 25.99 26.63
$\min_{K}e_{K_{y}}$ 2.375 2.312 2.112 1.965 1.807 1.620
$\max_{K}e_{K_{p}}$ 9.837 18.03 21.49 23.92 24.56 23.98
$\min_{K}e_{K_{p}}$ 1.406 1.294 1.284 1.333 1.222 1.093
Table 1. $e_{K_{y}}$ and $e_{K_{p}}$ of Example 4. Take $m=1,h_{u}=h.$
Next we show the image of $e_{K_{u}}$ for $m=1$. Here we plot $\min\{e_{K_{u}},2\}$ instead of $e_{K_{u}}$, for the reason that the maximum value of $e_{K_{u}}$ is too large and always exceeds 1000. We can see that the values of $e_{K_{u}}$ in the region $\{u^{*}>0\}$ become near to 1 with the refinement of $\mathcal{T}_{h},$ indicating that $\eta_{K_{u}}$ is efficient. The borderline between two regions $\{e_{K_{u}}>0\}$ and $\{e_{K_{u}}=0\}$ tends to the curve $\{u_{d}-p^{*}=0\}.$
Figure 12. Exact solution $u^{*}$ in Example 4.
Figure 13. $e_{K_{u}}$ of Example 4. Take $h_{u}=h=1/8.$
Figure 14. $e_{K_{u}}$ of Example 4. Take $h_{u}=h=1/32.$
Figure 15. $e_{K_{u}}$ of Example 4. Take $h_{u}=h=1/128.$
Example 5. We will present the results of $e_{K_{y}},e_{K_{p}}$ and $e_{K_{u}}$ for the problem in Example 2. We can see that the maximum and minimum values of $e_{K_{y}}$ and $e_{K_{p}}$ show no significant variation with the refinement of $\mathcal{T}_{h}$.
$h$ 1/4 1/8 1/16 1/32 1/64 1/128
$\max_{K}e_{K_{y}}$ 9.651 18.10 20.88 24.28 26.00 26.60
$\min_{K}e_{K_{y}}$ 2.376 2.313 2.111 1.964 1.807 1.620
$\max_{K}e_{K_{p}}$ 9.841 18.03 21.50 23.92 24.56 23.99
$\min_{K}e_{K_{p}}$ 1.407 1.294 1.285 1.333 1.222 1.093
Table 2. $e_{K_{y}}$ and $e_{K_{p}}$ of Example 5. Take $m=1,h_{u}=h.$
Next we show the results of $e_{K_{u}}$ for $m=1$. Consistent with Example 4, we plot $\min\{e_{K_{u}},4\}$ to avoid the influence of extreme values in $e_{K_{u}}$. In this case, our restriction is $U_{ad}=\{u\in L^{2}(\Omega_{u})\mid 0\leq u\leq 1\ \mathrm{a.e.\ in}\ \Omega_{u}\}$. Two clear borderlines are evident in the results, corresponding to the conditions $\{u_{d}-s(p^{*})=0\}$ and $\{u_{d}-s(p^{*})=1\}$. This is illustrated in Figures 17 and 19. Outside these boundaries, $e_{K_{u}}$ remains under 4.
Figure 16. Exact solution $u^{*}$ in Example 5.
Figure 17. $e_{K_{u}}$ of Example 5. Take $h_{u}=h=1/8.$
Figure 18. $e_{K_{u}}$ of Example 5. Take $h_{u}=h=1/32.$
Figure 19. $e_{K_{u}}$ of Example 5. Take $h_{u}=h=1/128.$
6. Conclusion
This work applies the reconstructed discontinuous approximation method to distributed elliptic optimal control problems with a strongly convex cost functional. We establish well-posedness of the discrete problem and derive both a priori and a posteriori error estimates in the $L^{2}$-norm and the energy norms. The discrete system is solved via a projected gradient descent algorithm with linear convergence rate. Numerical results substantiate the theoretical estimates and demonstrate the method’s practical effectiveness.
Acknowledgments
The research of this project is partially funded by National Natural Science Foundation of China (No. 12288101 and 12401399). The computational resources are partially supported by High-performance Computing Platform of Peking University.

### v2

@@
 Abstract.
-In this paper, we present and analyze an internal penalty discontinuous Galerkin method for the distributed elliptic optimal control problems.
+In this paper, we present and analyze an interior penalty discontinuous Galerkin method for the distributed elliptic optimal control problems.
 It is based on a reconstructed discontinuous approximation which admits arbitrarily high-order approximation space with only one unknown per element.
@@
 Our main contributions are twofold: (1) the derivation of both a priori and a posteriori error estimates of the $L^{2}$-norm and the energy norms, and (2) the implementation of an efficiently solvable discrete system, which is solved via a linearly convergent projected gradient descent method.
-Numerical experiments are provided to verify the convergence order in a priori estimate and the efficiency of a posteriori error estimate.
+Numerical experiments are provided to verify the convergence order in a priori error estimate and the efficiency of a posteriori error estimate.
 keywords: distributed elliptic control; discontinuous Galerkin method; reconstructed discontinuous approximation
@@
 2.1. Model Problem
-Let $\Omega_{u}\subset\mathbb{R}^{m}$ and $\Omega\subset\mathbb{R}^{d}(m,d=2,3)$ be convex polygonal or polyhedral domains.
+Let $\Omega_{u},\Omega\subset\mathbb{R}^{d}(d=2,3)$ be convex polygonal or polyhedral domains.
 Take the control space $U=L^{2}(\Omega_{u}),$ the state space $V=H^{2}(\Omega)$ and the admissible set $U_{ad}\subset U$ which is closed and convex.
@@
 $\lambda\|\xi\|^{2}\leq\xi^{T}A(x)\xi\leq\Lambda\|\xi\|^{2}\quad\forall\xi\in\mathbb{R}^{d},\ x\in\Omega.$
-In addition, we assume $f\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega),B:L^{2}(\Omega_{u})\rightarrow L^{2}(\Omega)$ is a bounded linear operator.
-The state equation of problem (2.1) can be divided into two independent parts as
-$\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{1})=f+Bu\quad\mathrm{in}\ \Omega,\\ &y_{1}=0\quad\mathrm{on}\ \partial\Omega,\end{aligned}\right.\quad\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{2})=0\quad\mathrm{in}\ \Omega,\\ &y_{2}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
-Notice that $y_{2}$ does not rely on $u$, we denote ${g_{1}}(y_{1}):=g(y_{1}+y_{2})$. Define a bilinear form $a:H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)\rightarrow\mathbb{R}$ as
-$a(y_{1},w):=(A\nabla y_{1},\nabla w)_{L^{2}(\Omega)}.$
-Denote the optimal solution to (2.1) by $(y^{\ast},u^{\ast})$, then we have the state equation
-(2.2) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*}\quad\mathrm{in}\ \Omega,\\ &y^{\ast}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
-and the relation $y_{1}^{\ast}=y^{\ast}-y_{2}$. Define the adjoint variable $p^{\ast}$ such that
-(2.3) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast})\quad\mathrm{in}\ \Omega,\quad\;\\ &p^{\ast}=0\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
-Consider the Lagrange function
-$L(y_{1}(u),u):={g_{1}}(y_{1}(u))+j(u)-a(y_{1}(u),p^{\ast})+(f+Bu,p^{\ast})\quad\forall u\in U_{ad}.$
-Here we use (,)($\cdot\,$$\cdot\,$) to denote the $L^{2}$ inner-product $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Notice that $L(y_{1}(u),u)=g_{1}(y_{1}(u))+j(u),\forall u\in U_{ad},$ we have
-$L(y_{1}(u^{\ast}+t(v-u^{\ast})),u^{\ast}+t(v-u^{\ast}))\geq L(y_{1}^{\ast},u^{\ast}),\quad\forall v\in U_{ad},\ t\in[0,1].$
-Differentiate the above inequality at $t=0$ in the direction $v-u^{*}.$ We have
-$(g_{1}^{\prime}(y_{1}^{*})y_{1}^{\prime}(u^{*})+j^{\prime}(u^{*}))(v-u^{*})-a(y_{1}^{\prime}(u^{*})(v-u^{*}),p^{*})+(B(v-u^{*}),p^{*})\geq 0.$
-Since
-$a(w,p^{*})=(g^{\prime}(y^{\ast}),w)=(g_{1}^{\prime}(y_{1}^{\ast}),w),\quad\forall w\in H_{0}^{1}(\Omega),$
-we take $w=y_{1}^{\prime}(u^{*})(v-u^{*})\in H_{0}^{1}(\Omega)$ and derive
-$(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
-So far we have derived the optimality conditions of problem (2.1) as follows.
-(2.4) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*},\quad y^{\ast}|_{\partial\Omega}=\phi,\\ &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast}),\quad p^{\ast}|_{\partial\Omega}=0,\\ &(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.\end{aligned}\right$.
-It is well-known from [23] that there exists a unique solution ($y^{*},u^{*}$) to problem (2.1), and the optimality condition (2.4) is a necessary and sufficient condition of the exact solution ($y^{*},u^{*}$).
+In addition, we assume $f\in L^{2}(\Omega),\phi\in H^{\frac{3}{2}}(\partial\Omega),B:L^{2}(\Omega_{u})\rightarrow L^{2}(\Omega)$ is a bounded linear operator. We further assume that
+Assumption 1.
+The functionals $j(u)$ and $g(y)$ can be written in the form
+$j(u)=\int_{\Omega_{u}}\tilde{j}(u)\mathrm{d}x,\quad g(y)=\int_{\Omega}\tilde{g}(y)\mathrm{d}x,$
+which means that
+$j^{\prime}(u)(v)=\int_{\Omega_{u}}\tilde{j}^{\prime}(u)v\mathrm{d}x,\quad g^{\prime}(y)(w)=\int_{\Omega}\tilde{g}^{\prime}(y)w\mathrm{d}x,$
+for $\forall u,v\in U,y,w\in Y.$
+[…]
 Assumption 2.
@@
 Here we denote $\|u-\tilde{u}\|_{L^{2}}:=\|u-\tilde{u}\|_{L^{2}(\Omega_{u})}$ and $\|y-\tilde{y}\|_{L^{2}}:=\|y-\tilde{y}\|_{L^{2}(\Omega)}.$ We will use these notations in the following part if there is no confusion.
-Assumption 2.
-The functionals $j(u)$ and $g(y)$ can be written in the form
-$j(u)=\int_{\Omega_{u}}\tilde{j}(u)\mathrm{d}x,\quad g(y)=\int_{\Omega}\tilde{g}(y)\mathrm{d}x,$
-which means that
-$(j^{\prime}(u),v)=\int_{\Omega_{u}}\tilde{j}^{\prime}(u)v\mathrm{d}x,\quad(g^{\prime}(y),w)=\int_{\Omega}\tilde{g}^{\prime}(y)w\mathrm{d}x,$
-for $\forall u,v\in U,y,w\in Y.$
-[…]
+Define a bilinear form $a:H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)\rightarrow\mathbb{R}$ as
+$a(y,w):=(A\nabla y,\nabla w)_{L^{2}(\Omega)}.$
+To derive the optimality condition of problem (2.1), we separate the state equation into two independent parts as
+$\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{1})=f+Bu\quad\mathrm{in}\ \Omega,\\ &y_{1}=0\quad\mathrm{on}\ \partial\Omega,\end{aligned}\right.\quad\left\{\begin{aligned} &-\nabla\cdot(A\nabla y_{2})=0\quad\mathrm{in}\ \Omega,\\ &y_{2}=\phi\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
+Notice that $y_{2}$ does not rely on $u$, we denote ${g_{1}}(y_{1}(u)):=g(y_{1}(u)+y_{2})$. Denote the optimal solution to (2.1) by $(y^{\ast},u^{\ast})$, then
+(2.2) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*}\quad\mathrm{in}\ \Omega,\\ &y^{\ast}=\phi\quad\mathrm{on}\ \partial\Omega,\end{aligned}\right$.
+and $y_{1}^{\ast}=y^{\ast}-y_{2}$. Define the adjoint variable $p^{\ast}$ such that
+(2.3) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast})\quad\mathrm{in}\ \Omega,\quad\;\\ &p^{\ast}=0\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
+Define the cost functional $J(y_{1}(u),u):={g_{1}}(y_{1}(u))+j(u),$ we have
+$J(y_{1}(u^{\ast}+t(v-u^{\ast})),u^{\ast}+t(v-u^{\ast}))\geq J(y_{1}^{\ast},u^{\ast}),\quad\forall v\in U_{ad},\ t\in[0,1].$
+Differentiate the above inequality at $t=0$ in the direction $v-u^{*},$ then
+$(g_{1}^{\prime}(y_{1}^{*})y_{1}^{\prime}(u^{*})+j^{\prime}(u^{*}))(v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
+Notice that $a(y_{1}(v),p^{*})=(f+Bv,p^{*})$ for $\forall v\in U_{ad}$, we have
+$a(y_{1}^{\prime}(u^{*})(v-u^{*}),p^{*})-(B(v-u^{*}),p^{*})=0,\quad\forall v\in U_{ad}.$
+Here we use (,)($\cdot\,$$\cdot\,$) to denote the $L^{2}$ inner-product $(\,\cdot\,,\,\cdot\,)_{L^{2}}$. Since
+$a(w,p^{*})=(g^{\prime}(y^{\ast}),w)=(g_{1}^{\prime}(y_{1}^{\ast}),w),\quad\forall w\in H_{0}^{1}(\Omega),$
+we take $w=y_{1}^{\prime}(u^{*})(v-u^{*})\in H_{0}^{1}(\Omega)$ and derive
+$(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.$
+So far we have derived the optimality conditions of problem (2.1) as follows.
+(2.4) $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y^{\ast})=f+Bu^{*},\quad y^{\ast}|_{\partial\Omega}=\phi,\\ &-\nabla\cdot(A\nabla p^{\ast})=g^{\prime}(y^{\ast}),\quad p^{\ast}|_{\partial\Omega}=0,\\ &(j^{\prime}(u^{*})+B^{*}p^{*},v-u^{*})\geq 0,\quad\forall v\in U_{ad}.\end{aligned}\right$.
+It is well-known from [23] that there exists a unique solution ($y^{*},u^{*}$) to problem (2.1), and the optimality condition (2.4) is a necessary and sufficient condition of the exact solution ($y^{*},u^{*}$).
 2.2. The Reconstruction Operator
@@
 $K_{1}:=\sup\limits_{K\in\mathcal{T}_{h}}\sup\limits_{e\in\mathcal{E}_{K}}\frac{h_{K}}{h_{e}}.$
-Here $\mathcal{E}_{K}=\{e\in\mathcal{E}_{h}\mid e\subset\partial K\}.$ So it does not change under uniform refinement of $\mathcal{T}_{h}.$
+Here $\mathcal{E}_{K}=\{e\in\mathcal{E}_{h}\mid e\subset\partial K\}.$ So it is invariant under uniform refinement of $\mathcal{T}_{h}.$
 We can extend the domain of $a_{h},l_{h}$ to $V_{h}$ without confusion, namely
@@
 In addition, for the exact solution $y$ to (2.6), we know by direct calculation that
-$a_{h}(y,w_{h}^{m})=l_{h}(w_{h}^{m}),\quad\forall w_{h}^{m}\in V_{h}^{m},$
-which leads to the consistency of $a_{h}$
 (2.9) $a_{h}(y,w_{h})=l_{h}(w_{h}),\quad\forall w_{h}\in V_{h}.$
@@
 Proof.
-Let $(y_{h}^{m*},u_{h}^{*})$ be the solution to (2.13). The Lagrange function
-$L(y_{h}^{m}(u_{h}),u_{h}):={g}(y_{h}^{m}(u_{h}))+j(u_{h})-(g^{\prime}(y_{h}^{m*}),y_{h}^{m}(u_{h})-F_{h}(f+Bu_{h},\phi)).$
+Let $(y_{h}^{m*},u_{h}^{*})$ be the solution to (2.13). Notice that
+$J(y_{h}^{m}(u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast})),u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast}))\geq J(y_{h}^{m*},u_{h}^{*}),\quad\forall v_{h}\in U_{adh},\ t\in[0,1].$
+Differentiate the above inequality at $t=0$ in the direction $v_{h}-u_{h}^{*},$ we have
+$(\tilde{g}^{\prime}(y_{h}^{m*})y_{h}^{m\prime}(u_{h}^{*})+\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})\geq 0,\quad\forall v_{h}\in U_{adh}.$
 […]
 $\left\{\begin{aligned} &-\nabla\cdot(A\nabla y)=w\quad\mathrm{in}\ \Omega,\\ &y=q\quad\mathrm{on}\ \partial\Omega.\end{aligned}\right$.
-We have $L(y_{h}^{m}(u_{h}),u_{h})=g(y_{h}^{m}(u_{h}))+j(u_{h}),\forall u_{h}\in U_{adh}.$ So
-$L(y_{h}^{m}(u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast})),u_{h}^{\ast}+t(v_{h}-u_{h}^{\ast}))\geq L(y_{h}^{m*},u_{h}^{*}),\quad\forall v_{h}\in U_{adh},\ t\in[0,1].$
-Differentiate the above inequality at $t=0$ in the direction $v_{h}-u_{h}^{*},$ we have
-$(\tilde{g}^{\prime}(y_{h}^{m*})y_{h}^{m\prime}(u_{h}^{*})+\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})$
-- $(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m\prime}(u_{h}^{*})(v_{h}-u_{h}^{*})-F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
-Here we use Assumption 2. Then
-$(\tilde{j}^{\prime}(u_{h}^{*}),v_{h}-u_{h}^{*})+(\tilde{g}^{\prime}(y_{h}^{m*}),F_{h}(B(v_{h}-u_{h}^{*}),0))\geq 0.$
-For $\forall w\in L^{2}(\Omega),$ we have
+Using the equality
+$(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m}(v_{h})-F_{h}(f+Bv_{h},\phi))=0,\quad\forall v_{h}\in U_{adh},$
+we deduce that
+$(\tilde{g}^{\prime}(y_{h}^{m*}),y_{h}^{m\prime}(u_{h}^{*})(v_{h}-u_{h}^{*})-F_{h}(B(v_{h}-u_{h}^{*}),0))=0,\quad\forall v_{h}\in U_{adh}.$
+In addition, for $\forall w\in L^{2}(\Omega),$
 $(\tilde{g}^{\prime}(y_{h}^{m*}),F_{h}(w,0))=l_{h}[\tilde{g}^{\prime}(y_{h}^{m*}),0](F_{h}(w,0))$
@@
 $=$ $l_{h}[w,0](F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0))=(F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0),w).$
+Take $w=B(v_{h}-u_{h}^{*})$ to derive
+$(\tilde{j}^{\prime}(u_{h}^{*})+B^{*}F_{h}(\tilde{g}^{\prime}(y_{h}^{m*}),0),v_{h}-u_{h}^{*})\geq 0,\quad\forall v_{h}\in U_{adh}.$
 […]
@@
 Construct the intermediate states as follows.
-$-\nabla\cdot(A\nabla Sy)=f+Bu_{h}\quad Sy|_{\partial\Omega}=\phi,$
+$-\nabla\cdot(A\nabla Sy)=f+Bu_{h},\quad Sy|_{\partial\Omega}=\phi,$
 $-\nabla\cdot(A\nabla Sp)=g^{\prime}(y_{h}^{m}),\quad Sp|_{\partial\Omega}=0,$
@@
 In this case, our restriction is $U_{ad}=\{u\in L^{2}(\Omega_{u})\mid 0\leq u\leq 1\ \mathrm{a.e.\ in}\ \Omega_{u}\}$.
-Two clear borderlines are evident in the results, corresponding to the conditions $\{u_{d}-s(p^{*})=0\}$ and $\{u_{d}-s(p^{*})=1\}$.
+Two borderlines are evident in the results, corresponding to the conditions $\{u_{d}-s(p^{*})=0\}$ and $\{u_{d}-s(p^{*})=1\}$.
 This is illustrated in Figures 17 and 19.